### v3

# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

Abstract
In binary decision-making, individuals often choose either the rare or the common action. In the framework of evolutionary game theory, the best-response update rule can be used to model this dichotomy. Those who prefer the common action are called coordinators, and those who prefer the rare one are called anticoordinators. A finite mixed population of the two types may undergo perpetual fluctuations, the characterization of which appears to be challenging. It is particularly unknown whether the fluctuations persist as population size grows. To fill this gap, we approximate the discrete finite population dynamics of coordinators and anticoordinators with the associated mean dynamics in the form of differential inclusions. We show that the family of the state sequences of the discrete dynamics for increasing population sizes forms a generalized stochastic approximation process for the differential inclusion. On the other hand, we show that the differential inclusions always converge to an equilibrium. This implies that the reported perpetual fluctuations in the finite discrete dynamics of coordinators and anticoordinators almost surely vanish with population size. The results encourage to first analyze the often simpler continuous-time mean dynamics of the discrete population dynamics as the continuous-time dynamics partly reveal the asymptotic behavior of the discrete dynamics.
Index Terms:
decision-making dynamics, best-response, mean-dynamics, evolutionary game theory
I Introduction
Humans face a variety of repetitive decision-making problems, such as whether to follow the fashion trends, get a flu shot, or sign a petition [1, 2]. In such two-option decision-making problems, individuals are often either conformists, also known as coordinators, those who adopt a decision whenever it has already been adopted by a particular population proportion of decision-makers, or nonconformists, also known as anticoordinators, those who go for a decision whenever it has been rejected by a particular population proportion [3, 4].
In the framework of evolutionary game theory, the so-called best-response update rule captures the decision-making processes of coordinators and anticoordinators by assuming linear payoff functions. An increasing payoff function with respect to the commonality of the decisions makes the best-response update rule equivalent to the linear threshold model, where the individuals are coordinators with possibly unique thresholds [5]. A decreasing payoff function, on the other hand, models the anticoordinators.
Analyzing these decision-making processes requires accounting for each individual’s actions, and consequently, a large body of research studied the steady state behavior of finite discrete populations of best-responders. The problem has been investigated in in various settings, ranging from well-mixed [6] to structured populations [7, 8, 4], and from time-invariant to time-varying thresholds [9]. Semi-tensor products were also utilized to analyze networked evolutionary games [10, 11] and logical dynamic games [12, 13] .
A finite population of heterogeneous coordinators, who differ in perceiving the commonality, admits equilibria [14] and equilibrates under asynchronous activation sequence [15]. A finite population of heterogeneous anticoordinators either equilibrates or converges to a two-state cycle [16, 17]. As for finite populations of coordinating and anticoordinating agents, it was shown that the dynamics may undergo perpetual fluctuations [18, 17, 19]. The exact form and the conditions for the existence of the fluctuations appear to be an open problem.
Researchers get around the challenges arising from the analysis of finite population dynamics by exploiting the associated deterministic mean dynamics [20, 21], where the population is assumed to be infinitely large. In this regard, most studies investigated the replicator dynamics which are the mean dynamics of the other popular update rule, imitation [22, 23, 24, 25]. Fewer studies considered the best-response mean dynamics, which are differential inclusions [26, 27, 28, 29, 30]. Some studies approximated the evolution of decisions in structured populations with ordinary differential equations [31, 32, 33].
Although analysis of the mean dynamics is generally more straightforward compared to that of finite populations, the potential discrepancy between these two may question the validity of such an approximation. Hence, much effort has been devoted to connecting the behavior of finite populations with the associated mean dynamics as the population size grows [34, 35, 36, 37]. The results on infinite horizon behavior connect the Birkhoff center of the mean dynamics with the support of the stationary measures of the Markov chain corresponding to the finite population’s dynamics [35, 38].
How can these results, which link the mean dynamics to finite population dynamics, be applied to the heterogeneous populations of coordinators and anticoordinators? To the best of our knowledge, no study investigated the asymptotic behavior of infinite heterogeneous populations of exclusively coordinators, exclusively anticoordinators, or a mixture of both. Even if there were such results, could they shed light on the discrete population dynamics?
We provide the answer in this paper, namely by using the deterministic differential inclusions as an approximation tool to reveal the asymptotic behavior of the discrete population dynamics as the population size approaches infinity. First, through an intuitive example we show that the perpetual fluctuations do not scale as the population size grows. Second, we write the population dynamics as a Markov chain and obtain the associated continuous-time population dynamics–an upper semicontinuous differential inclusion. Third, we show that the collection of population dynamics Markov chains define generalized stochastic approximation processes for the continuous-time population dynamics [38]. This allows the use of the results in [38] to approximate the asymptotic behavior of the discrete population dynamics by finding the Birkhoff center of the continuous-time population dynamics, which we do in the last part of the paper.
Our contribution is fourfold: i) We show that in the continuous-time population dynamics, the exclusively heterogeneous population of coordinators may admit two types of equilibrium points, clean-cut and ruffled coordinator-driven, where clean-cut equilibria are asymptotically stable and ruffled are unstable–Lemma 2 and Theorem 2. On the contrary, the exclusively heterogeneous population of anticoordinators admit only one globally asymptotically stable equilibrium point that could be either clean-cut or ruffled anti-coordinator driven–Lemma 2 and 1. These results are in line with the existing results on the steady-state behavior of finite heterogeneous populations of exclusively coordinators [15] and exclusively anticoordinators [16]. ii) We show that in the continuous-time population dynamics, the mixed heterogeneous population of coordinators and anticoordinators may admit all three types of equilibria, clean-cut, ruffled anticoordinator-driven, and ruffled coordinator-driven, where clean-cut and ruffled anticoordinator-driven equilibria are asymptotically stable, the other type is unstable, and the dynamics always converge to one of the equilibria–Lemma 2 and Theorem 2. iii) Consequently, we show that the reported perpetual fluctuations in the finite discrete dynamics of coordinators and anticoordinators [19] almost surely vanish with the population size –Theorem 3 and 4, which are the main result of this paper. iv) Out of the analysis of the continuous-time population dynamics, insightful hints on the asymptotic behavior of the finite populations were obtained, such as having the same equilibria and similar long-term behavior. Hence, in general for other population dynamics, it may prove useful to study the behavior of the continuous-time population dynamics, which is usually more straightforward, prior to the precise analysis of the finite population dynamics.
Notations
The set of real, nonnegative real, integer, nonnegative integer, and positive integer numbers are respectively shown by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, $\mathbb{Z}$, $\mathbb{Z}_{\geq 0}$, and $\mathbb{N}$. With an abuse of notation, for positive integers $n$ and $m$, define $\frac{1}{m}\mathbb{Z}^{n}=\{\frac{1}{m}\bm{y}\mid\bm{y}\in\mathbb{Z}^{n}\}$. For each $i\in\mathbb{N}$, $[i]$ denotes $\{1,2,\ldots,i\}$. Boldface letters refer to vectors. The notation $\bm{1}$ refers to a vector of appropriate size with all elements equal to $1$. The $i^{\text{th}}$ element of the vector $\mathbf{q}$ is indicated by ${q}_{i}$. The calligraphic font $\mathcal{X}$ represents a set. A sequence of variables $x_{0},x_{1},x_{2},\ldots$ is represented by $\langle x_{k}\rangle{{{}_{k=0}^{\infty}}}$. The Euclidean norm of a vector $\bm{x}$ is denoted by $|\bm{x}|$. By the notations $[a_{1},b_{1}]-[a_{2},b_{2}]$ and $[a_{1},b_{1}]-{c}$ we mean $[a_{1}-a_{2},b_{1}-b_{2}]$ and $[a_{1}-c,b_{1}-c]$, respectively. The notation $\text{cl}(\mathcal{X})$ denotes the closure of the set $\mathcal{X}$. A set-valued map ${\bm{\mathcal{V}}}(\bm{x})$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$ is denoted by notation ${\bm{\mathcal{V}}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$.
II Problem Formulation
We consider a well-mixed population of $\mathsf{N}$ agents, labeled by $1,2,\ldots,\mathsf{N}$, choosing repeatedly between two strategies $\mathtt{A}$ and $\mathtt{B}$ over a discrete time sequence $t\in\frac{1}{\mathsf{N}}\mathbb{Z}_{\geq 0}$. The time sequence is indexed by $k$ where $k=\mathsf{N}t$ (i.e., the index of time $\frac{k}{\mathsf{N}}$ is $k$). Each agent $i\in[\mathsf{N}]$ has a threshold $\tau(i)\in(0,1)$ and is either a coordinator (conformist) or an anticoordinator (a nonconformist). A coordinator $i$ (resp. an anticoordinator $i$) tends to choose strategy $\mathtt{A}$ whenever the population proportion of strategy-$\mathtt{A}$ players, excluding herself, denoted $x^{\mathsf{N}}_{{-i}}\in[0,1]$, does not fall short of (resp. does not exceed) her threshold, i.e., $x^{\mathsf{N}}_{{-i}}\geq\tau(i)$ (resp. $\leq\tau(i)$).
At each time index $k$, exactly one agent chosen uniformly at random becomes active and receives the chance to switch to her preferred strategy. At time index $k+1$, the strategy of agent $i$ active at time index $k$ will be
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}_{{-i}}(k)\geq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$ (1)
if she is a coordinator and otherwise, if she is an anticoordinator,
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}_{{-i}}(k)\leq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$ (2)
where $\mathtt{s}_{i}$ is the strategy of agent $i$.
Coordinators (resp. anticoordinators) who have the same threshold build up a subpopulation, and there all altogether $\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$) subpopulations of coordinators (resp. anticoordinators). Coordinating (resp. anticoordinating) subpopulations are labeled in the ascending (resp. descending) order of their thresholds by $1,\ldots,\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$), that is, $\tau^{\prime}_{1}<\tau^{\prime}_{2}<\ldots<\tau^{\prime}_{\mathsf{p}^{\prime}}$ (resp. $\tau_{1}>\tau_{2}>\ldots>\tau_{\mathsf{p}}$), where $\tau^{\prime}_{p}$ (resp. $\tau_{p}$) is the threshold of the $p^{\text{th}}$ coordinating (resp. anticoordinating) subpopulation. So the heterogeneity of the population is captured by the distribution of population proportions over the total $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations as $\bm{\rho}=({\rho}_{1},\ldots,{\rho}_{\mathsf{p}},{\rho}^{\prime}_{\mathsf{p}^{\prime}},\ldots,{\rho}^{\prime}_{1})^{\top}$ where ${\rho}^{\prime}_{j}$ (resp. ${\rho}_{j}$) represents the ratio of the number of individuals in coordinating (resp. anticoordinating) subpopulation $j$ to the population size $\mathsf{N}$. For convenience of vector indexing, for $j\geq\mathsf{p}+1$, we define ${\rho}_{j}={\rho}^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-j}$. Note that there is no subpopulation of size zero, meaning that $\min_{j}\{\rho_{j}\}\geq\frac{1}{\mathsf{N}}$.
Remark 1.
According to the best-response update rule, the active agent chooses the strategy which maximizes her current payoff, i.e, $s_{i}(k+1)=\mathtt{A}$ if $u_{i}^{\mathtt{A}}(k)\geq u_{i}^{\mathtt{B}}(k)$ and $\mathtt{s}_{i}(k+1)=\mathtt{B}$, otherwise where $u_{i}^{\mathtt{j}}(k)$ is the payoff of agent $i$ for choosing strategy $\mathtt{j}\in\{\mathtt{A},\mathtt{B}\}$ at time index $k$. If the agents’ payoffs are a linear function of the population proportion of strategy-$\mathtt{A}$ players, update rules (1) and (2) are equivalent to the best-response update rule.
At the population level, the collective behavior of the agents in each subpopulation is of interest rather than that of each single agent per se. As a result, we define the population state as the distribution of $\mathtt{A}$-players over the $(\mathsf{p}+\mathsf{p}^{\prime})$ subpopulations at each index $k$ and denote it by
$\bm{x}^{\mathsf{N}}=(\mathit{x}^{\mathsf{N}}_{1},\ldots,\mathit{x}^{\mathsf{N}}_{{\mathsf{p}}},\mathit{x}^{\prime\mathsf{N}}_{\mathsf{p}^{\prime}},\ldots,\mathit{x}^{\prime\mathsf{N}}_{{1}})^{\top},$
where $\mathit{x}^{\prime\mathsf{N}}_{p}$ (resp. $\mathit{x}^{\mathsf{N}}_{p}$) represents the proportion of $\mathtt{A}$-players who belong to coordinating (resp. anticoordinating) subpopulation $p$. We again use the notation $x^{\mathsf{N}}_{p}=x^{\prime\mathsf{N}}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}$ for $p>\mathsf{p}$. By defining
$\bm{\mathcal{X}}_{ss}=\prod_{j=1}^{{\mathsf{p}}}[0,{\rho}_{j}]\times\prod_{j=1}^{{\mathsf{p}^{\prime}}}[0,\rho^{\prime}_{\mathsf{p}^{\prime}-j+1}],$
the resulting state space then equals $\bm{\mathcal{X}}_{ss}\cap{\frac{1}{\mathsf{N}}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$.
The evolution of the state defines the population dynamics and are governed by update rules (1), (2), and the activation sequence of the agents. More specifically, the activation sequence is generated by the sequence of random variables $\langle{A}_{k}\rangle{{{}_{k=0}^{\infty}}}$ where ${A}_{k}$ is the agent active at time index $k$ and takes values in $[\mathsf{N}]$ with the distribution $\mathbb{P}[{A}_{k}=i]=\frac{1}{\mathsf{N}}$. A realization of an activation sequence $\langle{A}_{k}\rangle_{k=0}^{\infty}$ is persistent [39], if for every agent $i$ and every $K>0$ there exists a time index $T>K$ such that $A_{T}=i$.
As shown in A3, the above definition of population dynamics is equivalent to the following compact version.
Definition 1.
The discrete best-response population dynamics are defined by the following discrete time stochastic equation for $k=0,1,\ldots$
$\bm{x}^{\mathsf{N}}(k+1)\!$ $=\!\bm{x}^{\mathsf{N}}(k)\!+\!\frac{1}{\mathsf{N}}\bigg{(}\!{S}_{k}-{s^{*}}\big{(}{{P}_{k}},\bm{x}^{\mathsf{N}}{(k)},{S_{k}}\big{)}\!\bigg{)}\mathbf{e}_{{P}_{k}},$ (3)
where ${P}_{k}$ and ${S}_{k}$ are scalar random variables with distributions $\mathbb{P}[{P}_{k}=p]=\rho_{p}$, $\mathbb{P}[{S}_{k}=1|{P}_{k}=p]=x^{\mathsf{N}}_{p}/\rho_{p}$, $\mathbb{P}[{S}_{k}=2|{P}_{k}=p]=1-x^{\mathsf{N}}_{p}/\rho_{p}$, $\mathbf{e}_{{P}_{k}}$ is ${P}_{k}$th standard basis vector in $\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$, and ${s^{*}}(p,\bm{x}^{\mathsf{N}},{s})$ is a function where ${s^{*}}(p,\bm{x}^{\mathsf{N}},{1})$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of an $\mathtt{A}$-playing agent and ${s^{*}}(p,\bm{x}^{\mathsf{N}},{2})$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of a $\mathtt{B}$-playing agent in subpopulation $p$ at population state $\bm{x}^{\mathsf{N}}$ :
${s^{*}}($ $p,\bm{x}^{\mathsf{N}},{S_{k}})$
$=$ $\begin{cases}{1}&\text{if }(x^{\mathsf{N}}\leq{\tau_{p}}\text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}\leq{\tau_{p}}{+\frac{1}{\mathsf{N}}}\text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=1})\\ &\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}{+\frac{1}{\mathsf{N}}}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=1}),\\ {2}&\text{if }(x^{\mathsf{N}}>{\tau_{p}}\ \text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}>{\tau_{p}}{+\frac{1}{\mathsf{N}}}\ \text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=1})\\ &\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}{+\frac{1}{\mathsf{N}}}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=1}),\end{cases}$
where $x^{\mathsf{N}}=\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{p}^{\mathsf{N}}$.
In Definition 1, the random variable ${P}_{k}$ is the subpopulation of the agent active at time index $k$, and the random variable ${S}_{k}$ equals $1$ (resp. $2$) if the strategy of the active agent at index $k$ equals $\mathtt{A}$ (resp. $\mathtt{B}$). Persistent realizations of the sequences of random variables $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$ refer to their realizations under a persistent activation sequence. Function $s^{*}(p,\bm{x}^{\mathsf{N}},\cdot)$ returns $1$ if $\mathtt{A}$ is the preferred strategy of an agent in the subpopulation $p$ when the population is at state $\bm{x}^{\mathsf{N}}$, and it returns $2$ otherwise.
Example 1.
In the context of public health, when a newly developed vaccine is introduced, some individuals choose to get vaccinated (i.e., strategy $\mathtt{A}$) as long as the vaccination coverage is below a certain threshold [40]. For them, this signals that not enough individuals are vaccinated, and they want to be immunized. Conversely, some individuals go for vaccination when a specific proportion of the population has already been vaccinated [41] as then they perceive the vaccine as safe or a societal norm. The first group can be thought of as anticoordinators, while the latter are coordinators. The perception of what proportion of the population needs to be vaccinated to be considered rare or common varies among individuals, resulting in a heterogeneous population.
What is the asymptotic behavior of the population dynamics as the population size approaches infinity? Does the population state reach an equilibrium point where no agent tends to switch her adopted strategy? Or do the population dynamics undergo perpetual fluctuations?
We define a fluctuation set as a non-empty subset of the state space that is (i) positively invariant, meaning that once the state enters the set it never leaves it under any activation sequence, (ii) it is minimal, meaning that it does not admit a positively invariant proper subset, implying that once the state enters it, every state in the set will be visited under a persistent activation sequence, and (iii) it is non-singleton, implying that it is not an equilibrium. In what follows, we provide the definition in terms of the dynamics in Definition 1.
Definition 2.
A fluctuation set for the dynamics (3) is a non-singleton set $\bm{\mathcal{Y}}^{\mathsf{N}}\subseteq\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$ such that if $\bm{x}^{\mathsf{N}}(k)\in\bm{\mathcal{Y}}^{\mathsf{N}}$, for some $k$, then (i) $\bm{x}^{\mathsf{N}}(k^{\prime})\in\bm{\mathcal{Y}}^{\mathsf{N}}$ for all $k^{\prime}>k$ and (ii) for every state $\bm{y}^{\mathsf{N}}\in\bm{\mathcal{Y}}^{\mathsf{N}}$ and every $K>0$ and under every persistent realizations of $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$ , there exists some $T>K$ such that $\bm{x}^{\mathsf{N}}(k+T)=\bm{y}^{\mathsf{N}}$. The amplitude of the perpetual fluctuation set $\bm{\mathcal{Y}}^{\mathsf{N}}$ is defined as $\max_{\bm{y}^{\mathsf{N}},\bm{z}^{\mathsf{N}}\in\bm{\mathcal{Y}}^{\mathsf{N}}}|\bm{1}^{\top}\bm{y}^{\mathsf{N}}-\bm{1}^{\top}\bm{z}^{\mathsf{N}}|.$ Starting from a given initial condition and under a specified realization of $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$, we say that the dynamics (3) undergo perpetual fluctuations if $\bm{x}^{\mathsf{N}}(k)\in\bm{\mathcal{Y}}^{\mathsf{N}}$ for some fluctuation set $\bm{\mathcal{Y}}^{\mathsf{N}}$ and finite $k\geq 0$.
The dynamics of a finite mixed population consisting of both coordinating and anticoordinating subpopulations may undergo perpetual fluctuations [19], the characterization of which appears to be challenging and remain unsolved. Could the asymptotic behavior of discrete populations be revealed by investigating the associated mean dynamics, as the population size approaches infinity?
Figure 1: The long-term fluctuations of the population proportion of strategy-$\mathtt{A}$ players for varying population sizes. The circles and crosses represent the maximum and minimum of the population proportion of $\mathtt{A}$-players, respectively. As the population size increases, these values get closer.
Example 2.
Consider a population of $\mathsf{N}$ agents stratified into one anticoordinating and five coordinating subpopulations. The thresholds and population proportions of the subpopulations are respectively as follows: $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau^{\prime}_{5}},{\tau^{\prime}_{4}},{\tau^{\prime}_{3}},{\tau^{\prime}_{2}},{\tau^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.885},{0.89},{0.604},,{0.481},{0.444},{0.21})}$ $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho^{\prime}_{5}},{\rho^{\prime}_{4}},{\rho^{\prime}_{3}},{\rho^{\prime}_{2}},{\rho^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}12/30},{3/30},{3/30},{3/30},{1/30},{8/30}})$. For each value of $\mathsf{N}=30,60,120,\ldots,3840$, the population state was simulated 100 times with different random activation sequences and initial conditions. For each value of $\mathsf{N}$, we considered $30\mathsf{N}$ steps. For each run, the maximum and minimum of the population proportion of strategy $\mathtt{A}$ players were recorded. These values for each population size $\mathsf{N}$ are depicted in Figure 1. As population size grows, the maximum and minimum of the proportion of strategy-$\mathtt{A}$ players get closer and eventually merge.
Example 2 gives us an intuition that the amplitude of fluctuations in the population proportion of $\mathtt{A}$-players decreases as population size grows. Does this observation extend to every mixed population of coordinators and anticoordinators as the population size approaches infinity i.e., $\mathsf{N}\rightarrow\infty$? (Fig 2). The available results in stochastic approximation theory provide the basis for investigating this question.
Equilibrates [15]Equilibrates [16]May Fluctuate [19] Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both ? Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both $t\ \rightarrow\ \infty$$t\ \rightarrow\ \infty$ Population Population
Figure 2: How does the asymptotic behavior of the finite population of interacting agents (3) evolve as the population size approaches infinity? An exclusive finite population of coordinators and an exclusive finite population of anticoordinators each equilibrates in the long-run, [15, 16]. In the latter case, for some specific configurations, the population may admit a two-state cycle. The proportion of $\mathtt{A}$-players in a mixed finite population may undergo perpetual fluctuations [19].
III Background
We present the following definitions laying the basis for introducing Theorem 1, which connects the asymptotic behavior of the finite populations and their associated mean dynamics. Please see the appendix for more details.
A differential inclusion is defined by
$\dot{\bm{x}}(t)\in\bm{\mathcal{V}}\big{(}\bm{x}(t)\big{)},$ (4)
where $\bm{\mathcal{V}}\subset\mathbb{R}^{n}$ is a set-valued map specifying a set of evolution at each point $\bm{x}\in\mathbb{R}^{n}$.
Definition 3 ([38]).
Let $\bm{\Phi}$ be the dynamical system induced by (4). The recurrent points of $\bm{\Phi}$ are defined as
[…]
Each $\bm{y}$ in Definition 3 is a solution of the differential inclusion, and the term $\bigcap_{t\geq 0}\text{cl}(\bm{y}[t,\infty])$ is the limit set of $\bm{y}$. In words, $\bm{x}_{0}$ is a recurrent point of $\bm{\Phi}$ if it is included in the limit set of some solution that starts from $\bm{x}_{0}$. Namely, a recurrent point generalizes the notion of an equilibrium point as, for example, the points on a limit cycle are recurrent points.
How to approximate the realizations of a collection of Markov chains $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ for a sequence of $\epsilon>0$ with solutions of a differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$? The following definition establishes the basis for doing so.
Definition 4.
[38] Consider the good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ over the convex and compact state space $\bm{\mathcal{X}}_{0}$. For a sequence of positive values of scalar $\epsilon$ approaching 00, let $\textbf{\em U}^{\epsilon}=\langle\textbf{\em U}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ be a sequence of $\mathbb{R}^{n}$-valued random variables and $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ be a family of set-valued maps on $\mathbb{R}^{n}$. We say that $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\epsilon>0}$ is a family of generalized stochastic approximation processes (or a GSAP) for the differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ if the following conditions are satisfied:
  1. 1.
    $\mathbf{X}^{\epsilon}_{k}\in{\bm{\mathcal{X}}_{0}}$ for all $k\geq 0$,
  2. 2.
    $\mathbf{X}^{\epsilon}_{k+1}-\mathbf{X}^{\epsilon}_{k}-\epsilon\textbf{\em U}^{\epsilon}_{k+1}\in\epsilon\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$,
  3. 3.
    for any $\delta>0$, there exists an $\epsilon_{0}>0$ such that for all $\epsilon\leq\epsilon_{0}$ and $\bm{x}\in\bm{\mathcal{X}}_{0}$,
    $\bm{\mathcal{V}}^{\epsilon}(\bm{x})\subset\{\bm{z}$ $\in\mathbb{R}^{n}\mid\exists\bm{y}:|\bm{x}-\bm{y}|<\delta,$ (5)
    $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta\},$
  4. 4.
    for all $T>0$ and for all $\alpha>0$,
    $\lim_{\epsilon\rightarrow 0}\mathbb{P}\left[\max_{k\leq\frac{T}{\epsilon}}\left|\textstyle\sum_{i=1}^{k}\epsilon\textbf{\em U}^{\epsilon}_{i}\right|>\alpha\mid\mathbf{X}^{\epsilon}_{0}=\bm{x}\right]=0$
    uniformly in $\bm{x}\in\bm{\mathcal{X}}_{0}$.
The GSAPs can be regarded as perturbed solutions of the differential inclusions. The first condition ensures that the perturbed solutions are within the same set as the unperturbed solutions are. The second condition defines $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ as a sequence of elements which are recursively updated. At each index $k$, the value of $\mathbf{X}^{\epsilon}_{k+1}$ is characterized based on those of $\mathbf{X}^{\epsilon}_{k}$, $\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$, and the random variable $\textbf{\em U}^{\epsilon}_{k+1}$. The third condition ensures that the collection $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ approaches to the vicinity of $\bm{\mathcal{V}}$ as $\epsilon$ approaches zero. More specifically, the third condition is indeed a limit condition, which requires that for an arbitrarily small $\delta>0$ there exists $\epsilon_{0}>0$ such that for each $\epsilon<\epsilon_{0}$ in the sequence $\langle\epsilon\rangle$, the values belonging to the set valued map $\bm{\mathcal{V}}^{\epsilon}(\bm{x})$ are within a $\delta$-neighborhood of the image of the $\delta$-neighborhood of $\bm{x}$. Well-behaved random variables, such as sub-Gaussian random variables and some classes of random variables with finite higher moments satisfy the last condition.
In [38], Roth and Sandholm studied the limiting stationary distributions of these GSAPs and showed that if for each $\epsilon>0$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle$ is a Markov chain, their stationary measures as $\epsilon$ approaches zero will be concentrated on $\mathrm{BC}(\bm{\Phi})$. The following theorem is a result of A1 and was implicitly used in the proof of [38, Theorem 4.11].
Theorem 1.
For a vanishing sequence $\langle\epsilon_{n}\rangle_{n=0}^{\infty}$ let $\langle\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{n}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon_{n}$, $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ is a Markov chain and ${\mu}^{\epsilon_{n}}$ is an invariant probability measure of $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}$. We then have $\lim_{n\to\infty}{\mu}^{\epsilon_{n}}(\bm{\mathcal{O}})=1$, where $\bm{\mathcal{O}}$ is any open set containing the Birkhoff center of the dynamical system $\bm{\Phi}$ induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
IV The link to the continuous-time dynamics
Our aim is to determine the asymptotic behavior of the population dynamics (3) as the population size $\mathsf{N}$ approaches infinity. Theorem 1 connects the asymptotic behavior of the GSAPs to the steady-state behavior of their associated differential inclusion. Hence, if one shows that, firstly, dynamics (3) define a Markov chain, and secondly, the collection of the Markov chains indexed by the population size is a GSAP for a good upper semicontinuous differential inclusion, then the results of Theorem 1 can be applied, meaning that the asymptotic behavior of the population dynamics can be revealed by analyzing the differential inclusion.
In view of (3), the next state of the population dynamics only depends on the current population state and the adopted strategy of and the subpopulation associated with the current active agent. As a result, the sequence of $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{{}_{k=0}^{\infty}}}$ is a realization of a Markov chain.
Proposition 1.
The sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{{}_{k=0}^{\infty}}}$ is a realization of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0)$ and transition probabilities
$\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}=$
$\begin{cases}({\rho}_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))&\text{if }\bm{y}^{\mathsf{N}}=\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=-\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ {1-}\displaystyle\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\big{(}x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=\bm{x}^{\mathsf{N}},\\ \hskip 27.0pt+(\rho_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))\big{)}\\ 0&\text{otherwise},\end{cases}$ (6)
where $\mathbf{e}_{p}$ is the $p^{\text{th}}$ standard basis vector in $\mathbb{R}^{(\mathsf{p}+\mathsf{p}^{\prime})}$.
For the sake of readability, the proofs of the results are provided in the appendix.
Definition 5.
The population dynamics Markov chain is the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0),$ and transition probabilities (6).
We increase the population size $\mathsf{N}$ in a way that the population structure, i.e., the population proportions of the subpopulations $\bm{\rho}$ remains unchanged. Hence, the sequence according to which the population size approaches infinity should satisfy $\mathsf{N}\bm{\rho}\in\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$. We denote the set of such valid population sizes by $\mathcal{N}$.
Now, to apply Theorem 1, we need to show that the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for a good upper semicontinuous differential inclusion and hence the support of the limit point of its invariant probability measures is determined by the steady-state behavior of that differential inclusion. We claim that $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the following differential inclusion:
Definition 6.
The continuous-time population dynamics are defined by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ where $\bm{\mathcal{V}}:\bm{\mathcal{X}}_{ss}\rightrightarrows\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$, for $p\leq\mathsf{p}$
${\mathcal{V}}_{p}(\bm{x})=\begin{cases}\{{\rho}_{p}-x_{p}\}&\text{if }x<\tau_{p},\\ [-x_{p},{\rho}_{p}-x_{p}]&\text{if }x=\tau_{p},\\ \{-x_{p}\}&\text{if }x>\tau_{p},\end{cases}$ (7a)
and for $p>\mathsf{p}$
$\mathcal{V}_{p}(\bm{x})=\begin{cases}\{-x_{p}\}&\text{if }x<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ [-x_{p},{\rho}_{p}-x_{p}]&\text{if }x=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ \{{\rho}_{p}-x_{p}\}&\text{if }x>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\end{cases}$ (7b)
where $x=\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{p}$.
The above dynamics are indeed the mean dynamics of the discrete population dynamics (3) [20]. The mean dynamic considers the decision-making process as a continuous process with an infinitely large population, and $\dot{x}_{p}$ describes the rate of change in the population proportion of strategy-$\mathtt{A}$ players in subpopulation $p$. This equals the expected inflow,$F_{p,\mathtt{A}}$, that is, the rate of change to strategy $\mathtt{A}$ minus the expected outflow, $F_{p,\mathtt{B}}$, that is, the rate of change to strategy $\mathtt{B}$:
$\dot{x}_{p}\in{\{}F_{p,\mathtt{A}}(\bm{x})-F_{p,\mathtt{B}}(\bm{x}){\}},$ (8)
where $F_{p,\mathtt{A}},F_{p,\mathtt{B}}:\bm{\mathcal{X}}_{ss}\rightarrow{[0,\rho_{p}]}$. If the population proportion of strategy-$\mathtt{A}$ players is greater (resp. less) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the inflow will be zero, and the outflow equals the current proportion of strategy-$\mathtt{A}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=-x_{p}$. On the contrary, if the population proportion of strategy-$\mathtt{A}$ players is less (resp. greater) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the outflow will be zero and the inflow equals the current proportion of strategy $\mathtt{B}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=\rho_{p}-x_{p}$. When the population proportion of strategy-$\mathtt{A}$ players equals the threshold of (anticoordinating or coordinating) subpopulation $p$, both strategies $\mathtt{A}$ and $\mathtt{B}$ are a legitimate choice should there be no tie-breaker, and some may choose strategy $\mathtt{A}$ while others may choose strategy $\mathtt{B}$. Therefore, the rate will be bounded by the above two cases, i.e., $-x_{p}$ and $\rho_{p}-x_{p}$, resulting in the interval $[-x_{p},\rho_{p}-x_{p}]$. If there is a tie-breaker, such as the set-up in this paper, then one of the strategies is preferred, resulting in one of the above two cases, which results in a selection from the defined differential inclusion. A selection from $\bm{\mathcal{V}}(\bm{x})$ is a singleton map $\bm{\nu}(\bm{x})\in\mathbb{R}^{n}$ which satisfies $\bm{\nu}(\bm{x})\in\bm{\mathcal{V}}(\bm{x})$ for all $\bm{x}$ in $\mathbb{R}^{n}$ [42]. That is why the mean dynamics of the discrete population dynamics (3) is sometimes considered as a selection from (7) [38]. It can be easily checked that the differential inclusion (7) satisfies the conditions of good upper semicontinuity.
Remark 2.
The equivalent equation to (8) in the context of best-response would be
$\dot{\bm{x}}\in\bm{\mathrm{Br}}(\bm{x})-\bm{x}$
where $\bm{\mathrm{Br}}:\bm{\mathcal{X}}_{ss}{\rightrightarrows}{\bm{\mathcal{X}}_{ss}}$ is the best-response correspondence where $\mathrm{Br}_{p}(\bm{x})$ returns that population proportion of strategy-$\mathtt{A}$ players of subpopulation $p$ maximizing their utility defined in the game-theoretic context of the linear threshold models (see Remark 1).
Lemma 1.
The collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for (7).
Based on Lemma 1, the population dynamics (3) define GSAPs for the differential inclusion (7). Hence, in view of Theorem 1 the support of the limit point of the stationary distributions of (6), as $\mathsf{N}\rightarrow\infty$, is contained in the Birkhoff center of $\bm{\Phi}$ induced by (7). In the next section, we determine the Birkhoff center of $\bm{\Phi}$.
Remark 3.
The set-up in (1), (2), and (3) is based on a tie-breaking rule [43] where the agents prefer strategy $\mathtt{A}$ when the population proportion of $\mathtt{A}$-players, excluding themselves, equals the agent’s threshold, i.e., $x^{\mathsf{N}}_{-i}=\tau(i)$ for every agent $i$. However, for other tie-breaking rules, such as when strategy $\mathtt{B}$ is preferred, or the agents uniformly randomly choose either of the strategies, Lemma 1 remains valid–A1. The reason is that a different tie-breaker will only change $s^{*}(\cdot,\cdot,\cdot)$ in Definition 1 and consequently the selection from (7) when $x=\tau_{p}$ where $p$ is the type of agent $i$. Accordingly, the dynamical behavior of the resulting selection is already captured by (7). Therefore, investigating the asymptotic behavior of (7) does reveal that of the discrete population dynamics (3) with such tie-breaking rules as population size approaches infinity.
V The analysis of the continuous-time dynamics
To analyze the continuous-time dynamics, the equilibrium points of (7) should be characterized. The following subsection provides an intuition for how analyzing the evolution of $x(t)$ in Definition 6, i.e., the population proportion of strategy-$\mathtt{A}$ players, helps that of continuous-time population dynamics.
V-A Intuition and example
Example 3.
Consider a mixed population of one anticoordinating and two coordinating subpopulations. The thresholds and the population proportions of the subpopulations are respectively as follows: $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau_{2}^{\prime}},{\tau_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.85},{0.75},{0.35})$ and $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho_{2}^{\prime}},{\rho_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.6},{0.3},{0.1})$. The continuous-time population dynamics are $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$, where $\bm{\mathcal{V}}=[\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}]^{\top}$ and
${\mathcal{V}}_{1}(\bm{x})=\begin{cases}\{0.6-x_{1}\}&\text{if }x<0.85,\\ [-x_{1},0.6-x_{1}]&\text{if }x=0.85,\\ \{-x_{1}(t)\}&\text{if }x>0.85,\end{cases}$ (9a)
$\mathcal{V}_{2}(\bm{x})=\begin{cases}\{-x_{2}\}&\text{if }x<0.75,\\ [-x_{2},0.3-x_{2}]&\text{if}\quad x=0.75,\\ \{0.3-x_{2}\}&\text{if }x>0.75,\end{cases}$ (9ab)
$\mathcal{V}_{3}(\bm{x})=\begin{cases}\{-x_{3}\}&\text{if }x<0.35,\\ [-x_{3},0.1-x_{3}]&\text{if}\quad x=0.35,\\ \{0.1-x_{3}\}&\text{if }x>0.35.\end{cases}$ (9ac)
  • Motivation. To find the equilibrium points, we should find the state populations $\bm{x}^{*}$ at which $\bm{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$. The combinations of all possible cases for $\mathcal{V}_{1},\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ should be considered based on the value of $x$. So if we already knew the value of $x$ at equilibrium, we would only need to investigate the active case in $\mathcal{V}_{1}$, $\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ to find $x_{1},x_{2},x_{3}$, and in turn $\bm{x}$ at equilibrium. Hence, we obtain the dynamics $x(t)$ as follows.
  • The abstract dynamics. The thresholds divide the unit interval into four disjoint open intervals, i.e., $(0,0.35)$, $(0.35,0.75)$, $(0.75,0.85)$, and $(0.85,1)$. When $x\in(0,0.35)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and $\dot{\bm{x}}$ will be equal to $(0.6-x_{1},-x_{2},-x_{3})$. The derivative of $x$, will then be $\dot{x}=0.6-x$. A similar procedure can be applied to the other intervals, resulting in $\dot{x}\in\mathcal{X}(x)$, where
    $\mathcal{X}(x)=\begin{cases}\{0.6-x\}\quad&\text{if }x\in[0,0.35),\\ [0.6-x,0.7-x]\quad&\text{if }x=0.35,\\ \{0.7-x\}\quad&\text{if }x\in(0.35,0.75),\\ [0.7-x,1-x]\quad&\text{if }x=0.75,\\ \{1-x\}\quad&\text{if }x\in(0.75,0.85),\\ [0.4-x,1-x]\quad&\text{if }x=0.85,\\ \{0.4-x\}\quad&\text{if }x\in(0.85,1].\end{cases}$
  • Abstract dynamics’ equilibria. At equilibrium, we have $0\in\mathcal{X}(x^{*})$. So we can find the equilibria by investigating all intervals in (10). The first two intervals are equilibrium-free. However, the image of the interval $(0.35,0.75)$ under $\mathcal{X}$ is $(-0.05,0.35)$, which includes zero. Therefore, the abstract dynamics admit an equilibrium point in this interval, which turns out to be $x^{*}_{1}=0.7$ as $0\in\mathcal{X}(0.7)$. Following the same procedure, we find two additional equilibrium points $x^{*}_{2}=0.75$ and $x^{*}_{3}=0.85$ for the abstract dynamics.
  • Population dynamics’ equilibria. In view of (9a), the value $x^{*}_{1}$ equals the sum of $\rho_{1}$ and $\rho^{\prime}_{1}$, i.e., the population proportions of those subpopulations whose preferred strategy at $x^{*}_{1}$ is $\mathtt{A}$. Evaluating (9a) at $x=x^{*}_{1}$, we see that $\bm{x}^{*}_{1}=(0.6,0,0.1)$ is an equilibrium point for the population dynamics. The point $x^{*}_{2}$ equals the threshold of the second coordinating subpopulation, and the population state $\bm{x}^{*}_{2}=(0.6,0.05,0.1)$ is an equilibrium point for (9a). At $\bm{x}^{*}_{2}$, except for the second coordinating subpopulation, members of the same subpopulations adopt the same strategies, i.e., $x^{*}_{2,1}=\rho_{1}$, $x^{*}_{2,3}=\rho^{\prime}_{1}$. Finally, the point $x^{*}_{3}=0.85$ equals the anticoordinating subpopulation’s threshold, and $\bm{x}^{*}_{3}=(0.45,0.3,0.1)$ is an equilibrium for the population dynamics. At $\bm{x}^{*}_{3}$, except for the anticoordinators, members of the same subpopulations adopt the same strategies, i.e., $x^{*}_{3,2}=\rho^{\prime}_{2}$, $x^{*}_{3,3}=\rho^{\prime}_{1}$.
V-B The abstract dynamics.
Example 3 gives us an intuition that it is useful to focus on the evolution of the population proportion of $\mathtt{A}$-players, i.e., $x(t)$, which from now on we refer to as the abstract state as it is abstracting away the heterogeneity of the population. Define ${\eta}^{\prime}_{j}$ (resp. ${\eta}_{i}$) as the cumulative population proportion of coordinating (resp. anticoordinating) subpopulations, that is the population proportion of those having thresholds equal to or less (resp. greater) than $\tau^{\prime}_{j}$ (resp. $\tau_{i}$), i.e.,
${\eta}^{\prime}_{j}=\sum_{k=1}^{j}\rho^{\prime}_{k},\quad{\eta}_{i}=\sum_{k=1}^{i}\rho_{k},$
where $j\in[\mathsf{p}^{\prime}]$, $i\in[\mathsf{p}]$, and we define ${\eta}^{\prime}_{0}={\eta}_{0}={\eta}_{-1}={\eta}^{\prime}_{-1}=0$, ${\eta}^{\prime}_{\mathsf{p}^{\prime}+1}={\eta}^{\prime}_{\mathsf{p}^{\prime}},{\eta}_{\mathsf{p}+1}={\eta}_{\mathsf{p}}$, $\tau_{\mathsf{p}+1}=\tau^{\prime}_{0}=0$, and $\tau_{0}=\tau^{\prime}_{\mathsf{p}^{\prime}+1}=1$.
Assumption 1.
The thresholds of the subpopulations are unique and satisfy the following:
$\forall k\!\in[\mathsf{p}]\cup\{0\}\forall l\in[\mathsf{p}^{\prime}]\cup\{0\}({\eta}_{k}+{\eta}^{\prime}_{l}\notin\!\{\tau_{1},\ldots,\tau_{\mathsf{p}},\tau^{\prime}_{1},\ldots,\tau^{\prime}_{\mathsf{p}^{\prime}}\}).$
In words, it is assumed that the sum of the cumulative population proportion of first $k$ anticoordinating subpopulations and first $l$ coordinating subpopulations does not match any of the thresholds for the $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations. We refer the reader to Remark 4 for consequences of this assumption.
Proposition 2.
Consider the continuous-time dynamics (7). Under 1, the evolution of the population proportion of $\mathtt{A}$-players $x(t)$ is governed by
$\dot{x}\in\mathcal{X}(x),$ (11)
[…]
where for $x=\tau_{i}$, $j$ equals $\max\{k|\tau^{\prime}_{k}<\tau_{i}\}$ and for $x=\tau_{j}^{\prime}$, $i$ equals $\max\{k|\tau^{\prime}_{j}<\tau_{k}\}$.
We refer to (11) as the abstract dynamics.
V-C Equilibrium points
Inspired by the definitions in [44] and based on Example 3, we claim that the equilibria of the population dynamics are either clean-cut or ruffled as defined in the following.
Definition 7.
A clean-cut population state, is a population state $\mathbf{c}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}],$ defined by
$\mathbf{c}^{ij}=({\rho}_{1},\ldots,{\rho}_{i},0,\ldots,0,{\rho}^{\prime}_{j},\ldots,{\rho}^{\prime}_{1}).$
The abstract state $x$ at $\mathbf{c}^{ij}$ is denoted by ${c}^{ij}={\eta}_{i}+{\eta}^{\prime}_{j}$. The state at which no coordinating (resp. anticoordinating) agents adopt strategy $\mathtt{A}$, is denoted by $\mathbf{c}^{i0}$ (resp. $\mathbf{c}^{0j}$).
In a clean-cut population state, all members of every subpopulation adopt the same strategy in a way that all coordinating (resp. anticoordinating) subpopulations with thresholds equal to or less (resp. greater) than that of some benchmark coordinating (resp. anticoordinating) subpopulation $j\in[\mathsf{p}^{\prime}]$ (resp. $i\in[\mathsf{p}]$) play strategy $\mathtt{A}$ and the remaining subpopulations play strategy $\mathtt{B}$.
Definition 8.
A ruffled anticoordinator-driven population state is a population state $\mathbf{a}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}],$ defined by
$\mathbf{a}^{ij}=\big{(}{\rho}_{1},\ldots,{\rho}_{i-1},\tau_{i}-({\eta}^{\prime}_{j}+{\eta}_{i-1}),0,\ldots,0,{\rho}^{\prime}_{j},\ldots,{\rho}^{\prime}_{1}\big{)},$
where $j$ equals $\max\{k\in[\mathsf{p}^{\prime}]\cup\{0\}|\tau^{\prime}_{k}<\tau_{i}\}$. The abstract state at $\mathbf{a}^{ij}$ is denoted by $a^{ij}=\tau_{i}$. For $j=0$, no coordinators play $\mathtt{A}$.
In a ruffled anticoordinator-driven population state, the population proportion of strategy-$\mathtt{A}$ players equals the threshold of some benchmark anticoordinating subpopulation $i$ such that all coordinating (resp. anticoordinating) subpopulations with a threshold less (resp. greater) than that of subpopulation $i$ play strategy $\mathtt{A}$.
Definition 9.
A ruffled coordinator-driven population state is a population state $\mathbf{o}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}]$, defined by
$\mathbf{o}^{ij}=\big{(}{\rho}_{1},\ldots,{\rho}_{i},0,\ldots,0,$ $\tau^{\prime}_{j}-({\eta}^{\prime}_{j-1}+{\eta}_{i}),{\rho}^{\prime}_{j-1}\ldots,{\rho}_{1}^{\prime}\big{)},$
where $i=\max\{k\in[\mathsf{p}]\cup\{0\}|\tau^{\prime}_{j}<\tau_{k}\}$. The abstract state at $\mathbf{o}^{ij}$ is denoted by $o^{ij}=\tau^{\prime}_{j}$. For $i=0$, no anticoordinators play $\mathtt{A}$.
In a ruffled coordinator-driven population state, the population proportion of strategy-$\mathtt{A}$ players equals the threshold of some benchmark coordinating subpopulation $j$ such that all coordinating (resp. anticoordinating) subpopulations with a threshold less (resp. greater) than that of subpopulation $j$ play strategy $\mathtt{A}$.
In Example 3, $\bm{x}^{*}_{1}$ (resp. $x^{*}_{1})$ is a clean-cut equilibrium point for (9a) (resp. (10)) and equals $\mathbf{c}^{11}$ (resp. ${c}^{11}$), and $\bm{x}^{*}_{3}$ (resp. $x^{*}_{3}$) is a ruffled anticoordinator-driven equilibrium point for (9a) (resp. (10)) and equals $\mathbf{a}^{12}$ (resp. ${a}^{12}$). Finally, $\bm{x}^{*}_{2}$ (resp. $x^{*}_{2}$) is a ruffled coordinator-driven equilibrium point for (9a) (resp. (10)) and equals $\mathbf{o}^{12}$ (resp. ${o}^{12}$). From now on, for the sake of readability, we drop the word “ruffled” and simply say “(anti)coordinator-driven.” The following lemma proves that the equilibria of the continuous-time population and abstract dynamics are one of these three population states.
Lemma 2.
The following hold under 1:
  1. 1.
    The equilibrium points of the dynamics (7) and (11) are either clean-cut, anticoordinator-driven or coordinator-driven population states.
  2. 2.
    ${c}^{ij}$ (resp. $\mathbf{c}^{ij}$) is an equilibrium of (11) (resp. (7)) iff
    $\max\{\tau_{i+1},\tau^{\prime}_{j}\}<{\eta}^{\prime}_{j}+{\eta}_{i}<\min\{\tau_{i},\tau^{\prime}_{j+1}\}.$ (12)
  3. 3.
    ${a}^{ij}$ (resp. $\mathbf{a}^{ij}$) is an equilibrium (11) (resp. (7)) iff
    $0<\tau_{i}-({\eta}^{\prime}_{j}+{\eta}_{i-1})<{\rho}_{i}.$ (13)
  4. 4.
    ${o}^{ij}$ (resp. $\mathbf{o}^{ij}$) is an equilibrium (11) (resp. (7)) iff
    ${\eta}_{i}+{\eta}^{\prime}_{j-1}<\tau^{\prime}_{j}<{\eta}_{i}+{\eta}^{\prime}_{j}.$ (14)
According to Lemma 2, for every equilibrium of the continuous-time population dynamics, e.g., $\mathbf{c}^{ij}$, there is exactly one associated abstract equilibrium point, i.e., $c^{ij}$.
V-D Global stability analysis
The abstract dynamics and, in turn, the continuous-time population dynamics may admit several clean-cut, anticoordinator-driven, or coordinator-driven equilibrium points. Denote the set of clean-cut, anticoordinator-driven, and coordinator-driven equilibrium points of the continuous-time population dynamics by $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$, respectively. When there is no equilibrium point of either type, the corresponding set will be empty. Assume that there are altogether $\mathtt{Q}\in\mathbb{N}$ equilibrium points of clean-cut or anticoordinator-driven type. Arrange the clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics (7) in the ascending order of their associated abstract equilibrium points. Hence, we have
$q^{*}_{1}<q^{*}_{2}<\ldots<q^{*}_{\mathtt{Q}},$
where $q^{*}_{k}$ is the abstract state at $\mathbf{q}^{*}_{k}\in\bm{\mathcal{Q}}^{a}\cup\bm{\mathcal{Q}}^{c}$ for $k\in[\mathtt{Q}]$. A set $\bm{\mathcal{M}}$ is attractive under differential inclusion (4) from a set $\bm{\mathcal{U}}$ if for each solution $\bm{x}(t)$ with $\bm{x}(0)\in\bm{\mathcal{U}}$ and each open $\epsilon$-neighborhood of $\bm{\mathcal{M}}$, there exists time $T>0$ such that $\bm{x}(t)$ falls in $\epsilon$-neighborhood of $\bm{\mathcal{M}}$ for all $t\geq T$ [45]. The union of all sets from which $\bm{\mathcal{M}}$ is attractive under differential inclusion $\bm{\mathcal{V}}$ is the basin of attraction of $\bm{\mathcal{M}}$ under $\bm{\mathcal{V}}$ [45].
According to A2, in the abstract dynamics, between every two anticoordinator-driven and/or clean-cut equilibria $q^{*}_{k}$ and $q^{*}_{k+1}$, $k\in[\mathtt{Q}-1]$, there exists exactly one coordinator-driven equilibrium, which we denote by $q^{*}_{k,k+1}$. The associated equilibrium in the continuous-time dynamics is denoted by $\mathbf{q}^{*}_{k,k+1}$. Define $q^{*}_{0,1}=0$ and $q^{*}_{\mathtt{Q},\mathtt{Q}+1}=1$. The following theorem reveals the global behavior of the continuous-time population dynamics.
Theorem 2.
Consider the continuous-time population dynamics (7) and its associated abstract dynamics (11). Under 1,
  1. 1.
    Each $\mathbf{q}^{*}_{k-1,k},k\in\{2,\ldots,\mathtt{Q}\},$ is unstable, and each $\mathbf{q}^{*}_{k},k\in[\mathtt{Q}]$, is asymptotically stable with the basin of attraction
    $\bm{\mathcal{A}}(\mathbf{q}^{*}_{k})=\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})\}.$
  2. 2.
    The limit set of every point in $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x={q}^{*}_{k,k+1}\}$ for $k\in[\mathtt{Q}-1]$ is ${\{}\bm{q}^{*}_{k},\bm{q}^{*}_{k+1},\bm{q}^{*}_{k,k+1}{\}}$.
Indeed it can be shown that the clean-cut equilibrium points are exponentially asymptotically stable and the anticoordinator-driven equilibrium points are finite-time stable.
Corollary 1.
Under 1, for the continuous-time population dynamics (7) and the associated abstract dynamics (11),
  1. 1.
    there exists at least one clean-cut or anticoordinator-driven equilibrium point;
  2. 2.
    if the dynamics admit only one equilibrium point, it is globally asymptotically stable.
V-E The Birkhoff center of the continuous-time population dynamics
The following proposition determines the Birkhoff center of the dynamical system induced by (7).
Proposition 3.
If Assumption 1 holds, the Birkhoff center of the dynamical system induced by the continuous-time population dynamics (7) is
$\mathrm{BC}_{\text{M}}=\bm{\mathcal{Q}}^{c}\cup\bm{\mathcal{Q}}^{a}\cup\bm{\mathcal{Q}}^{o}.$ (15)
The following corollaries are direct results of Proposition 3 and determine the Birkhoff center of the dynamical system induced by the continuous-time population dynamics capturing the behavior of a population of exclusively coordinators and exclusively anticoordinators, respectively.
Corollary 2.
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics of a population of exclusively coordinators is
$\mathrm{BC}_{\text{C}}$ $=\bm{\mathcal{Q}}^{c}\cup\bm{\mathcal{Q}}^{o}.$ (16)
As for the population of anticoordinators, the result would be $\mathrm{BC}_{\text{A}}=\bm{\mathcal{Q}}^{a}\cup\bm{\mathcal{Q}}^{c}$. However, in this case, these sets can be simplified as it can be shown that there is only one equilibrium (1) and it is either a clean-cut or ruffled anticoordinator-driven. We summarize this possibility into the following state:
$\mathbf{q}=({\rho}_{1},\ldots,{\rho}_{p-1},\min\{\tau_{p}-{\eta}_{p-1},{\rho}_{p}\},0,\ldots,0).$
In view of Lemma 2, the condition for the existence of this equilibrium would be ${\eta}_{p-1}<\tau_{p}$ and $\tau_{p+1}<{\eta}_{p}$.
Corollary 3.
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics describing the evolution of an exclusive population of $\mathsf{p}$ types of anticoordinators is
$\mathrm{BC}_{\text{A}}$ =$=\{\mathbf{q}\}.$ (17)
So far, the steady-state behavior of the continuous-time population dynamics (7) was determined. The following section determines the asymptotic behavior of the discrete population dynamics (3) as the population size approaches infinity.
VI The asymptotic behavior of the discrete population dynamics
The following theorem asserts the main result of this paper.
Theorem 3.
Consider the discrete population dynamics for a population size $\mathsf{N}$ (3). Let ${\mu}^{\frac{1}{\mathsf{N}}}$ be an invariant probability measure of the corresponding population dynamics Markov chain. Under 1, for every vanishing sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ we have $\lim_{\frac{1}{\mathsf{N}}\to 0}{\mu}^{\frac{1}{\mathsf{N}}}(\bm{\mathcal{O}})=1$ where $\bm{\mathcal{O}}$ is any open set containing
  • $\mathrm{BC}_{\text{M}}$ if the population is a mixture of coordinating and anticoordinating subpopulations;
  • $\mathrm{BC}_{\text{A}}$ if the population consists of anticoordinating subpopulations;
  • $\mathrm{BC}_{\text{C}}$ if the population consists of coordinating subpopulations.
In words, consider a population of coordinators and anticoordinators who update their decisions asynchronously based on (1) and (2), respectively. As the population size approaches infinity, the population state almost surely only visits the states close to the equilibria. Consequently, the amplitude of the possible perpetual fluctuations in the population proportion of $\mathtt{A}$-players will converge to zero. This results in the following corollary.
Corollary 4.
Consider the population dynamics (3). Under 1 and starting from a given initial condition, as the population size approaches infinity, the amplitude of the fluctuations in the population proportion of $\mathtt{A}$-players converges to zero with probability one.
Nevertheless, we can still have perpetual fluctuations (with nonzero probability), but the amplitude of the fluctuations will vanish with the population size almost surely. Roughly speaking, the state if does not equilibrate for “large” populations most of the time in the long run will spend “close” to the isolated members of the Birkhoff center.
Remark 4.
1 ensures that $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$ are distinct. Otherwise, the following possible cases may happen: (i) The threshold of benchmark anticoordinating subpopulation $i$ equals ${\eta}^{\prime}_{j}+{\eta}_{i-1}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{a}^{ij}$ and $\mathbf{c}^{(i-1)j}$ (resp. $\mathbf{c}^{ij}$) match and remain stable. (ii) The threshold of benchmark coordinating subpopulation $j$ equals ${\eta}^{\prime}_{j-1}+{\eta}_{i}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{o}^{ij}$ and $\mathbf{c}^{i(j-1)}$ (resp. $\mathbf{c}^{ij}$) match, resulting in an unstable equilibrium. (iii) The threshold of benchmark coordinating subpopulation $j$ equals that of benchmark anticoordinating subpopulation $i$. In this case, the point $x=\tau_{i}$ is an equilibrium for the abstract dynamics if $\tau_{i}\in[{\eta}_{i-1}+{\eta}_{j-1}^{\prime},{\eta}_{i}+{\eta}_{j}^{\prime}]$. Following the approach taken in the proof of Theorem 2, it can be shown that if $\tau_{i}\in({\eta}_{i-1}+{\eta}_{j}^{\prime},{\eta}_{i}+{\eta}_{j-1}^{\prime})$, it is asymptotically stable, and unstable otherwise. In this case a continuum of equilibria $\mathbf{z}^{ij}$ emerges for the semicontinouous population dynamics. This continumm is characterized by $\bm{\mathcal{Z}}=\{\bm{x}\in\bm{\mathcal{X}}_{ss}|\forall l\notin\{i,\mathsf{p}+\mathsf{p}^{\prime}-j+1\},z^{ij}_{l}=a^{ij}_{l},z^{ij}_{i}+z^{ij}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}=\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})\}$. In words, all members of those coordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}^{\prime}$ play strategy $\mathtt{A}$ (resp. $\mathtt{B}$), and all members of those anticoordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}$ play strategy $\mathtt{B}$ (resp. $\mathtt{A}$). The proportions of strategy $\mathtt{A}$ players in the two benchmark subpopulations add up to $\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})$. The Birkhoff center will then also contain $\bm{\mathcal{Z}}$. The population proportion of strategy-$\mathtt{A}$ players is constant at $\bm{\mathcal{Z}}$ and consequently, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players in the discrete population dynamics vanished with probability one.
Remark 5.
In some setups, such as in [15], agents include themselves when determining their preferred strategy. As detailed in Section VII-D, our main results hold for such update rules.
Revisiting Example 2.
Consider the associated continuous-time population dynamics with the finite population dynamics introduced in Example 2. Based on Lemma 2, the mean dynamics admit only one equilibrium point at $x=\tau_{1}$, which is an anticoordinator-driven equilibrium point. Therefore, according to 1 this equilibrium point is globally asymptotically stable. Figure 3 depicts the evolution of population proportion of strategy-$\mathtt{A}$ players merged with Figure 1. As seen, $x(t)$ approaches $\tau_{1}$. The chattering observed in the solid black line is due to the numerical errors.
Example 4.
Consider a population of size $\mathsf{N}$ consisting of four coordinating and three anticoordinating subpopulations. The population distribution over the subpopulations is $\bm{\rho}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{6/28},{8/28},{3/28},{3/28})$. The thresholds of the anticoordinating subpopulations are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.929}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{2}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6/7}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{3}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.357}$, and those of coordinators are ${\tau^{\prime}_{1}}={0.05}$, ${\tau^{\prime}_{2}}={0.321}$, ${\tau^{\prime}_{3}}={0.5}$, ${\tau^{\prime}_{4}}={0.643}$. When the population size $\mathsf{N}$ equals $28$ or $56$ (which is the next larger valid size), the population proportion of $\mathtt{A}$-players either reaches $x^{\mathsf{N}}=11/28$ and remains there or fluctuates around $x^{\mathsf{N}}=6/7$. However, for larger populations, the fluctuation disappears.
Continuous-time population dynamics. In view of Lemma 2, the continuous-time population dynamics associated with the described discrete population dynamics admit three equilibrium points; one clean-cut equilibrium point $\mathbf{c}^{22}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{0},{0},{3/28},{3/28})$ with the corresponding abstract state $c^{22}$ equals $11/28$, one anticoordinator-driven equilibrium point $\mathbf{a}^{24}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{6/28},{8/28},{3/28},{3/28})$ with the corresponding abstract state $a^{24}$ equal to $6/7$, and one coordinator-driven equilibrium point $\mathbf{o}^{23}$ at $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3/28},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{0},{3/28},{3/28},{3/28})$ with the corresponding abstract state $o^{23}$ equal to the threshold of the third coordinating subpopulation, $0.5$. According to Theorem 2, except for a set of initial conditions of measure zero, the population state converges to either of $\mathbf{c}^{22}$ or $\mathbf{a}^{24}$, depending on the initial condition–Figure 4. As for the discrete population dynamics, the observed fluctuations are around ${a}^{24}$. In addition, the equilibrium state at $x^{\mathsf{N}}=11/28$ preserved as the population size grows and is indeed equal to ${c}^{22}$.
Remark 6.
Several studies have analyzed the population dynamics by studying one-dimensional dynamics. In [5], the discrete-time evolution of a heterogeneous population of coordinators was studied by analyzing the evolution of the population proportion of strategy-$\mathtt{A}$ players. A similar approach was used in [19, Section 9] for a heterogeneous population of coordinators and anticoordinators under a synchronous updating rule, where all agents are active at every time step. In [31, 32], scalar continuous-time dynamics instead serve as approximations of the evolution of the population proportion of strategy-$\mathtt{A}$ players in structured populations. In this paper, however, we were interested in studying the evolution of the proportion of strategy-$\mathtt{A}$ players within each subpopulation. Hence, a one-dimensional dynamic would not reveal the behavior of each sub-population. The dimension should be at least as many as the number of sub-populations, i.e., $\mathsf{p}+\mathsf{p}^{\prime}$. Moreover, the evolution of the total proportion of $\mathtt{A}$-players, i.e., $x^{\mathsf{N}}$, is not a Markov chain because the change from $x^{\mathsf{N}}$ to $x^{\mathsf{N}}+1/\mathsf{N}$ (or to $x^{\mathsf{N}}-1/\mathsf{N}$) depends on the proportions of $\mathtt{A}$-players in subpopulations. Nevertheless, the expected change of $x^{\mathsf{N}}$ can be calculated, and that is how we provided the abstract dynamics, which facilitated the subsequent analysis of the $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional population dynamics.
Figure 3: The associated abstract state with the discrete population dynamics described in Example 2 approaches $\tau_{1}$. The black solid line represents the evolution of the abstract state over time. The circles and crosses represent the upper and lower bounds of the invariant sets for different population sizes.
Figure 4: Example 4. The solution of the continuous-time population dynamics for two different initial conditions. The upper panel shows the evolution of the population state which converges to clean-cut equilibrium point $\mathbf{c}^{22}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,3,0},{0,0,3,3})/28$ and the lower one depicts the evolution of the population state converging to the anticoordinator-driven equilibrium point $\mathbf{a}^{24}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,2,0},{6,8,3,3})/28$, where the corresponding population proportion of $\mathtt{A}$-players equals $\tau_{2}=6/7$. The observed non-smoothness in the evolution of some subpopulations is due to the change in their preferred strategy. For example, in the upper panel, at time zero, the abstract state is below $\tau^{\prime}_{2}$ (resp. $\tau_{3}$) resulting in $\dot{x}^{\prime}_{2}<0$ ($\dot{x}_{3}>0$). As abstract state increases and exceeds $\tau^{\prime}_{2}$ (resp. $\tau_{3}$), the preferred strategy of coordinating (resp. anticoordinating) subpopulation $2$ (resp. $3$) changes to $\mathtt{A}$ (resp. $\mathtt{B}$) yielding $\dot{x}^{\prime}_{2}>0$ (resp. $\dot{x}_{3}<0$).
VII Concluding Remarks
We analyzed the asymptotic behavior of a well-mixed finite heterogeneous population of agents repeatedly choosing between two strategies when the population size approaches infinity. We utilized the available results linking the stationary measures of the Markov chain corresponding to this discrete population dynamics with the steady-state behavior of the associated continuous-time mean dynamics. We then showed that the fluctuations in the finite mixed population of coordinators and anticoordinators likely do not persist with the population size (Figure 2).
The convergence result of the continuous-time population dynamics for exclusive populations of anticoordinators, i.e., 3, matches that of the discrete dynamics [16] where the population state was proven to always reach the clean-cut equilibrium $\mathbf{q}$ or reach the set of two ruffled states (that are not possible here due to the continuity of the state space of the continuous-time population dynamics). Similarly, the convergence result of the continuous-time population dynamics for exclusive populations of coordinators, i.e., 2, matches that of the discrete dynamics [15] where the population state was proven to always equilibrate. Moreover, the equilibria of the discrete and continuous-time dynamics match.
In addition, the stability of clean-cut equilibrium points of continuous-time population dynamics matches that of the discrete dynamics [46]. However, the stability of the ruffled types has not been investigated in the discrete population.
In case of the fluctuations, in all of the examples, we observed that the interval defined by the minimum and maximum of the ratio of strategy $\mathtt{A}$-players in the same population and for a fixed initial condition and different activation sequence contains an anticoordinator-driven equilibrium point of the abstract dynamics. Moreover, as the population size grows, the interval shrinks. It remains yet unknown whether this observation holds true always for every population and whether ruffled anticoordinator-driven equilibria are necessary for the existence of minimal positively invariant sets in the finite discrete populations. The difficulty partly arises from the inherent challenge of characterizing the positively invariant sets and amplitude of the fluctuations. Although [19, Corollary 6] provides bounds on the fluctuations, it still requires the benchmark subpopulations to be known, which appears to be an unsolved complicated problem. The result of this paper holds true in a well-mixed population. Whether a similar result can be obtained for structured populations, such as k-regular networks, is an open question. In this regard, approaches similar to those in [32] might work.
continuous-time population dynamics Stochastic approximation theory Coordinating Anticoordinating Mixture of both $t\ \rightarrow\ \infty$EquilibratesEquilibratesEquilibratesThe fluctuations vanish Population EquilibratesEquilibratesMay Fluctuate Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both $t\ \rightarrow\ \infty$$t\ \rightarrow\ \infty$ Population Population
Figure 2: [Revisit]-The connection between the asymptotic behavior of finite populations of interacting agents evolving based on (3) and their associated continuous-time population dynamics (7) based on Stochastic approximation theory. As population size approaches infinity, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players vanishes with probability $1$.
Appendix-Basic Definitions
[…]
A Caratheodory solution of (4) on $[0,t_{1}]\subset\mathbb{R}_{\geq 0}$ is an absolutely continuous map $\bm{x}^{o}:[0,t_{1}]\rightarrow\mathbb{R}^{n}$ such that for almost all $t\in[0,t_{1}]$, $\dot{\bm{x}}^{o}(t)\in\bm{\mathcal{V}}(\bm{x}^{o}(t))$ [51]. From now on, by a solution of (4), we mean its Caratheodory solution. Robotic manipulators interacting with objects, high-speed supercavitation vehicles, and power systems can be well modeled and analyzed within the framework of differential inclusions [51, 52, 53]. A point $\bm{x}^{*}\in\mathbb{R}^{n}$ that satisfies ${\mathbf{0}}\in\bm{\mathcal{V}}(\bm{x}^{*})$ is an equilibrium of (4) [51]. The set-valued dynamical system induced by the differential inclusion (4) is denoted by $\bm{\Phi}(\bm{x}_{0})$ and defined as the set of all solutions of (4) with initial condition $\bm{x}_{0}$ [38].
Example A 1.
Consider the scalar differential inclusion (4) with $x\in\mathbb{R}$ and
$\mathcal{V}\big{(}x(t)\big{)}=\begin{cases}\{-1-x(t)\}&\text{if }x(t)<0,\\ [-1,1]&\text{if }x(t)=0,\\ \{1-x(t)\}&\text{if }x(t)>0.\end{cases}$
For any positive initial condition $x(0)$, the solution of (4) equals $x(t)=1+e^{-t}(x(0)-1)$. For any negative initial condition $x(0)$, the solution of (4) equals $x(t)=-1+e^{-t}(1+x(0))$. For the initial condition $x(0)=0$, however, the solution $x(t)$ is not unique, see Figure 3. The set valued dynamical system $\bm{\Phi}$ for initial condition $x(0)=0$ contains $\{0,e^{-t}-1,1-e^{-t}\}$, and for other initial conditions will be a singleton. As for equilibrium points of (4), we need to find the points $x^{*}$ at which $0\in\mathcal{V}(x^{*})$, yielding $-1$, 00, and $1$.
Figure 3: 1. The upper panel represents the set-valued map $\mathcal{V}(x)$. The lower panel represents three solutions of $\dot{x}(t)\in\mathcal{V}(x)$ with 00 as the initial state, i.e., $0,e^{-t}-1,$ and $1-e^{-t}$.
A set-valued map $\bm{\mathcal{V}}:\bm{\mathcal{X}}\rightrightarrows\bm{\mathcal{Y}}$ is called upper semicontinuous if for every sequence $\langle\bm{x}_{n}\rangle_{{n=0}}^{{\infty}}$, $\bm{x}_{n}\in\bm{\mathcal{X}}$, with limit point $\bm{x}_{0}\in\bm{\mathcal{X}}$ and every sequence $\langle\bm{y}_{n}\rangle_{{n=0}}^{{\infty}}$, $\bm{y}_{n}\in\bm{\mathcal{V}}(\bm{x}_{n})$, with limit point $\bm{y}_{0}$, we have $\bm{y}_{0}\in\bm{\mathcal{V}}(\bm{x}_{0})$. When an upper semicontinuous map is equipped with three additional properties of nonemptiness, convexity, and boundedness, the map is good upper semicontinuous. As for the following scalar set-valued map
$\mathcal{V}(x(t))=\begin{cases}\{-1-x(t)\}&\text{if }x(t)<0,\\ [0,1]&\text{if }x(t)=0,\\ \{1-x(t)\}&\text{if }x(t)>0,\end{cases}$
consider the sequence $\langle x_{n}\rangle_{n}$, where $x_{n}=-0.5/(n+1)$, for $n=0,1,\ldots$. The sequence is associated with $\langle y_{n}\rangle_{n}$, where $y_{n}=-1+0.5/(n+1)$. The limit point of the sequence $\langle x_{n}\rangle_{n}$ is $x_{0}=0$ and that of the sequence $\langle y_{n}\rangle_{n}$ is $y_{0}=-1$. However, $y_{0}\notin\mathcal{V}(x_{0})$. Hence, the map is not upper semicontinuous, because it is not so at $x=0$. A sequence $\langle\epsilon_{k}\rangle_{{k=0}}^{{\infty}}$ of positive constants that converges to zero is called vanishing.
Theorem A1.
[38, Theorem 3.5 and Corollary 3.9] For a vanishing sequence of $\epsilon>0$, let $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\epsilon>0}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}$ is a Markov chain and let ${\mu}^{\epsilon}$ be an invariant probability measure of $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$. Let ${\mu}$ be a limit point of $\langle{\mu}^{\epsilon}\rangle_{\epsilon>0}$ in the topology of weak convergence. Then the support of ${\mu}$ is contained in $\mathrm{BC}(\bm{\Phi})$, where $\bm{\Phi}$ is the dynamical system induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
Proofs
VII-A Proof of Theorem 1
Proof.
Assume by contradiction that $\lim_{k\to\infty}{{\mu}}^{\epsilon_{k}}(\bm{\mathcal{{O}}})\not\to 1$. This implies that for some positive constant $\delta<1$, there exists a subsequence $\langle\alpha_{k}\rangle_{k=1}^{\infty}=\alpha_{1},\alpha_{2},\ldots$ of $\langle\epsilon_{k}\rangle_{k=0}^{\infty}$ such that ${{\mu}}^{\alpha_{k}}(\bm{\mathcal{{O}}})<\delta$ for each $k$. As the state space $\bm{\mathcal{X}}_{0}$ is compact, the family of probability measures defined on the space $\bm{\mathcal{X}}_{0}$ is tight [50]. Then, in view of Prokhorov’s theorem, every sequence of the probability measures, including $\langle{\mu}^{\alpha}\rangle$ has a weak limit point, which is a probability measure ${\mu^{\prime}}$. This implies that there exists a subsequence $\langle\beta_{k}\rangle_{k=1}^{\infty}=\beta_{1},\beta_{2},\ldots$ of $\langle\alpha_{k}\rangle_{k=1}^{\infty}$ such that $\lim\inf_{k\to\infty}{{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})\geq{{\mu^{\prime}}}(\bm{\mathcal{{O}}})$. On the other hand, in view of A1, the support of every limit point is contained in the Birkhoff center $\mathrm{BC}(\bm{\Phi})$, i.e., ${\mu}^{\prime}(\bm{\mathcal{O}})=1$. This means that there exists some $k^{*}$ such that for all $k>k^{*}$ ${{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})>\delta$, a contradiction. ∎
VII-B Proof of 1
Proof.
To prove, it suffices to show that ${\mathbb{P}[\mathbf{X}_{k+1}=\bm{x}^{\mathsf{N}}(k+1)|\mathbf{X}_{k}=\bm{x}^{\mathsf{N}}(k)]=\text{Pr}_{\bm{x}^{\mathsf{N}}(k),\bm{x}^{\mathsf{N}}(k+1)}}$. That is to show that the probability the population dynamics (3) reach the state $\bm{y}^{\mathsf{N}}$ at index $k+1$ from the state $\bm{x}^{\mathsf{N}}$ at index $k$ equals the transition probability $\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}$ in the Markov chain. This is straightforward, because at any state $\bm{x}^{\mathsf{N}}$ either the active agent switches to $\mathtt{A}$ or to $\mathtt{B}$ or sticks to her current strategy and no other case is possible. This summarizes the possible four cases in (6). In the first case, for example, the probability of an agent that belongs to the subpopulation $p$ switches her strategy from $\mathtt{B}$ to $\mathtt{A}$ equals the probability of drawing such an agent ${\rho}_{p}-x^{\mathsf{N}}_{p}$ if strategy $\mathtt{A}$ is the preferred strategy of subpopulation $p$’s agents with current strategy $\mathtt{B}$, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}},{2})=1}$, and equals zero otherwise, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}},{2})=2}$. ∎
VII-C Proof of Lemma 1
First, we introduce the following lemma, which facilitates the proof of Lemma 1.
Lemma A1.
Under 1, let $d$ denote the minimum distance between thresholds, i.e., $d=\min_{i,j\in[\mathsf{p}+\mathsf{p}^{\prime}]}|\tau_{i}-\tau_{j}|.$ Then, for any population size $\mathsf{N}$ satisfying $\mathsf{N}d>1$ and any $\bm{x}\in\bm{\mathcal{X}}_{ss}$, the sum $\bm{1}^{\top}\bm{x}$ belongs to at most one interval of the form $[\tau_{i},\tau_{i}+\frac{1}{\mathsf{N}})$ with $i\in[\mathsf{p}+\mathsf{p}^{\prime}]$.
Proof.
We assume, on the contrary, that there is also a threshold $\tau_{j}$ where $\bm{1}^{\top}\bm{x}\in[\tau_{j},\tau_{j}+\frac{1}{\mathsf{N}})$ and that $\tau_{i}\neq\tau_{j}$. Without loss of generality, let $\tau_{j}>\tau_{i}$ which results in $\tau_{j}<x<\tau_{i}+1/\mathsf{N}$ and, in turn, $\tau_{j}<\tau_{i}+1/\mathsf{N}$. However, according to the assumption $1/\mathsf{N}<d$, and we reach a contradiction. ∎
The proof of Lemma 1 is inspired by the steps taken in [38, Example 4.1].
Proof.
By considering the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence, it suffices to show that there exist $\langle\bm{\mathcal{V}}^{\frac{1}{\mathsf{N}}}\rangle_{{\frac{1}{\mathsf{N}}}}$ and $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ such that the conditions in Definition 4 are met. The first condition of Definition 4 can be easily verified. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$. Since $k=\mathsf{N}t$, there are $\mathsf{N}$ transitions per unit time resulting in $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, which is the expected increment in the subpopulation $p$, equals the sum of the multiplication of each possible change in the subpopulation $p$ with its probability. In view of (6), there are two possible changes, i.e., $\frac{1}{\mathsf{N}}$ step increment or $\frac{1}{\mathsf{N}}$ step reduction, which results in ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\mathsf{N}\big{(}\frac{1}{\mathsf{N}}(\rho_{p}-x^{\mathsf{N}}_{p})(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-\frac{1}{\mathsf{N}}(x^{\mathsf{N}}_{p}(s^{*}(p,\bm{x}^{\mathsf{N}},{1})-1))\big{)}$ and consequently, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-x^{\mathsf{N}}_{p}{+x^{\mathsf{N}}_{p}\left(s^{*}(p,\bm{x}^{\mathsf{N}},2)-s^{*}(p,\bm{x}^{\mathsf{N}},1)\right)}$. The second condition of Definition 4 is satisfied by $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}=$ $\smash{\mathsf{N}(\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}])}$ resulting in $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-{\frac{1}{\mathsf{N}}}\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}={\frac{1}{\mathsf{N}}}\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$. Now, we investigate Condition 3. To show that Condition 3 is satisfied, for a given $\delta$, we need to provide an $\epsilon_{0}$ satisfying eq. 5. We propose an $\epsilon_{0}$ which satisfies $\epsilon_{0}<\min\{d,\delta\}.$ Then, given A1, for any $\bm{x}\in\bm{\mathcal{X}}_{ss}$ there exists at most one threshold, say $\tau_{l}$, $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, that satisfies $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\frac{1}{\mathsf{N}})$, with $\frac{1}{\mathsf{N}}\leq\epsilon_{0}.$ In what follows, we show that Condition 3 holds for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ that satisfy $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$; the other values of $\bm{x}$ can be handled similarly. Let $\varepsilon_{x}=\bm{1}^{\top}\bm{x}-\tau_{l}$. Then given $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$, we have $\varepsilon_{x}<\epsilon_{0}$ and, in turn, $\varepsilon_{x}<\delta.$ Consider the point $\bm{y}$ whose all elements, except some element $j$, are equal to the corresponding elements of $\bm{x}$, and the $j^{\text{th}}$ element, $y_{j}$, equals $x_{j}-\varepsilon_{x}$. The term $|\bm{x}-\bm{y}|$ then equals $\varepsilon_{x}$, and hence the point $\bm{y}$, which satisfies $\bm{1}^{\top}\bm{y}=\tau_{l}$, belongs to the $\delta$-neighborhood of $\bm{x}$, i.e., $|\bm{x}-\bm{y}|<\delta$. For all $p\neq l$, we then have $\mathcal{V}_{p}({\bm{y}})=2(\rho_{p}-s^{*}(p,{\bm{y}},2))-{y}_{p}$ and for $p=l$, $\mathcal{V}_{p}=[-{y}_{p},\rho_{p}-{y}_{p}]$. On the other hand, the relation $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$ implies that ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ for $p=l$ equals $\rho_{l}-2x^{\frac{1}{\mathsf{N}}}_{l}$ (resp. 00) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation. For $p\neq l$, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ equals $2(\rho_{p}-s^{*}(p,\bm{x},2))-x_{p}$, and it can be easily shown that it is also equal to the unique member of $\mathcal{V}_{p}(\bm{x})$ defined in (7). By $\bm{z}=\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ and choosing $\bm{v}$ to be a member of $\bm{\mathcal{V}}({\bm{y}})$ with $v_{l}({\bm{y}})=\rho_{l}-2{y}_{l}$ (resp. $v_{l}({\bm{y}})=0$) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation, the term $|\bm{z}-\bm{v}|$ equals $|\bm{v}({\bm{y}})-\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})|=|{\bm{y}}-\bm{x}|$ which is less than $\delta$. This implies the satisfaction $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta$ and, in turn, Condition 3. As for the last condition, in view of Example 4.1 in [38], it can be shown that $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ is a martingale difference sequence. On the other hand, for all $k$, $|\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}|\leq\sqrt{\Sigma_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}(1+\rho_{l})^{2}}$ and hence $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}$ is uniformly bounded. Therefore, the conditions asserted in [38, Proposition 2.3] are satisfied and the satisfaction of the last condition is guaranteed. ∎
VII-D Proof of Lemma 1 when agents include themselves
As noted in Remark 3, in some set-ups, agents take into account their current strategies when they want to determine their proffered strategies. In this case, the update rule (1) changes to
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}(k)\geq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$
and the update rule (2) changes to
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}(k)\leq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$
where $x^{\mathsf{N}}$ is the population proportion of $\mathtt{A}$-players. Accordingly, the function $s^{*}(\cdot,\cdot,\cdot)$, defined in Definition 1 and used in Equation 6, changes to
${\hat{s}}(p,\bm{x}^{\mathsf{N}})=$
$\begin{cases}{1}&\text{if }(x^{\mathsf{N}}\leq{\tau_{p}}\text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}),\\ {2}&\text{if }(x^{\mathsf{N}}>{\tau_{p}}\ \text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}).\end{cases}$
If Lemma 1 remains true, then the main results of the papers are also valid. Condition 1 is already satisfied. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$, i.e., $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p$th element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, reads as ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-\hat{s}(p,\bm{x}^{\mathsf{N}}))-x^{\mathsf{N}}_{p}$. Accordingly, Conditions 2 and 4 can be easily verified. Regarding Condition 3, For each given $\delta$, we need to find an $\epsilon_{0}$ which satisfies (5). In what follows, we show that (5) is satisfied for any arbitrarily chosen $\epsilon_{0}$. We consider a selection from (7) denoted by $\dot{\bm{x}}=\bm{\nu}(\bm{x})$ where $\bm{\nu}$ is a $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional function and its $p$th element ${\nu}_{p}$ equals $\rho_{p}-x_{p}$ if $\mathtt{A}$ is the preferred strategy of subpopulation $p$ and equals $-x_{p}$ if the preferred strategy of subpopulation $p$ is $\mathtt{B}$. This results in ${\nu}_{p}(\bm{x})=\rho_{p}(2-\hat{s}(p,\bm{x}))-x_{p}$. The expected increment per unit of time, $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, is the same as the function $\bm{\nu}(\bm{x})$. As a result, equation (5) and, in turn, Condition 3 are satisfied by selecting $\bm{y}$ equal to $\bm{x}$, regardless of the value of $\epsilon_{0}$. This implies that Lemma 1 holds.
VII-E Proof of 2
Proof.
It holds that $x=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{k}$ and $0\leq x\leq 1$. Considering 1, the thresholds of the subpopulations are distinct, building $\mathsf{p}+\mathsf{p}^{\prime}+1$ disjoint open intervals in the unit interval, where their limit points are either 00, $1$, or the thresholds. Denote by $\mathcal{T}$ the union of the set $\{0,1\}$ with the set of all subpopulations’ thresholds. When $x\notin\mathcal{T}$, there exist $i\in\{0,1,\ldots,\mathsf{p}+1\}$ and $j\in\{0,1,\ldots,\mathsf{p}^{\prime}+1\}$ where $x\in(\tau^{\prime}_{j},\tau_{i})$ such that $\tau_{i+1}<x<\tau^{\prime}_{j+1}$. In view of (7), in this case, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and for all $0<l\leq i$ anticoordinating subpopulations, $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining anticoordinating subpopulations, i.e., $i<l\leq\mathsf{p}$, $\dot{x}_{l}=-x_{l}$. Similarly, for coordinating subpopulations, $\dot{x}_{l}=-x_{l}$ for all $\mathsf{p}<l\leq\mathsf{p}+\mathsf{p}^{\prime}-j+1$ and $\dot{x}_{l}=\rho_{l}-x_{l}$ for all $\mathsf{p}+\mathsf{p}^{\prime}-j+1<l\leq\mathsf{p}+\mathsf{p}^{\prime}$. By summing over the elements of $\dot{\bm{x}}$, we have $\dot{x}=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\dot{x}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}-x$. Therefore, for $x\in(\tau^{\prime}_{j},\tau_{i})$, $\mathcal{X}(x)=\{{\eta}_{i}+{\eta}^{\prime}_{j}-x\}$ which is the same as the last case of $\mathcal{X}$ in (11). When $x\in\mathcal{T}$, three cases may happen: $x=\tau^{\prime}_{j}$, $x=\tau_{i}$, or $x\in\{0,1\}$. When $x=\tau_{i}$, based on (7), except for $\mathcal{V}_{i}$, the elements of $\bm{\mathcal{V}}(\bm{x})$ are single-valued and contain either $\rho_{l}-x_{l}$ or $-x_{l}$. More specifically, for anticoordinating subpopulations, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ and $\dot{x}_{l}=-x_{l}$ for $i<l\leq\mathsf{p}$. As for coordinating subpopulations, assume that the set $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is not empty and let $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$. Then, for $l$ satisfying $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, $\dot{x}_{l}$ will be equal to $\rho_{l}-x_{l}$, and for the remaining subpopulations will be equal to $-x_{l}$. If $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is empty, then $\dot{x}_{l}$ for $l>\mathsf{p}$ will be equal to $-x_{l}$. Summing over the elements of $\dot{\bm{x}}$ except the $i^{\text{th}}$ element, yields ${\eta}_{i-1}+{\eta}^{\prime}_{j}-(x-x_{i})$. As for $l=i$, we have $\dot{x}_{i}\in[-x_{i},\rho_{i}-x_{i}]$ yielding $\dot{x}\in({\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x)$. Therefore, $\mathcal{X}(x)=[{\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the first case of $\mathcal{X}$ in (11). If $x=\tau^{\prime}_{j}$, a similar reasoning results in $\mathcal{X}(x)=[{\eta}_{i}+{\eta}^{\prime}_{j-1}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the second case of $\mathcal{X}$ in (11). Finally, when $x\in\{0,1\}$, a similar reasoning yields $\dot{x}={\eta}_{\mathsf{p}}-x$ if $x=0$ and $\dot{x}={\eta}^{\prime}_{\mathsf{p}^{\prime}}-x$ if $x=1$. These two are equivalent to the first two cases of $\mathcal{X}$ in (11) as $\tau_{\mathsf{p}+1}=\tau_{0}^{\prime}=0$ and $\tau^{\prime}_{\mathsf{p}^{\prime}+1}=\tau_{0}=1$. ∎
VII-F Proof of Lemma 2
Proof.
The proofs are first done for the abstract dynamics. We first prove parts 2, 3, and 4. Part 2) (sufficiency) Should (12) hold, it would follow that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$ and based on (11), $0\in\mathcal{X}(c^{ij})$, implying that $c^{ij}$ is an equilibrium of (11). (Necessity) Should $c^{ij}$ be an equilibrium of the abstract dynamics, then $0\in\mathcal{X}(c^{ij})$. One of the three cases of $\mathcal{X}$ in (11) must be active for $x=c^{ij}$. The first two cases of $\mathcal{X}$ cannot be active because of 1. So the third one must hold, implying that $c^{ij}\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$, which is equivalent to (12). Part 3) and Part 4). Following Part 2, the sufficiency and necessity can be concluded. Part 1) In view of 1, equilibrium points of clean-cut, anticoordinator-driven, and coordinator-driven types do not match for (11). Hence, an abstract state can at most be only one of these three types. We prove by contradiction that an equilibrium point of (11) should be either of these three types. Suppose $x^{*}$ is an equilibrium point for (11) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. At $x=x^{*}$ one of the cases of $\mathcal{X}$ in (11) should be active. If the first case is active, we have $x^{*}=\tau_{i}$ and as it is an equilibrium point $0\in\mathcal{X}(x^{*})$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<{\eta}_{i}+{\eta}^{\prime}_{j}$, which is equivalent to (13), but we assumed that $x^{*}$ is not an anticoordinator-driven equilibrium point. This is a contradiction and as a result the first case cannot be active. Similarly, for the remaining two cases, we reach a contradiction, and hence $x^{*}$ should be either clean-cut, anticoordinator-driven, or coordinator-driven. Now, we prove the lemma for the equilibrium points of the continuous-time population dynamics (7). The sufficiency and necessity of Part 2) can be easily shown. Part 3) (Sufficiency) Assume that (13) holds. By plugging $\mathbf{a}^{ij}$ into (7), we see that $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and hence $\mathbf{a}^{ij}$ is an equilibrium point for (7). (Necessity) Should $\mathbf{a}^{ij}$ be an equilibrium point for (7), $\mathbf{0}\in\bm{\mathcal{V}}(\mathbf{a}^{ij})$ and consequently ${0}\in\mathcal{V}_{i}(\mathbf{a}^{ij})$. This results in $0\in[-{a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$ and by plugging ${a}^{ij}_{i}=\tau_{i}-({\eta}_{i-1}+{\eta}^{\prime}_{j})$ into $[-{a}^{ij}_{i},\rho_{i}-{a}^{ij}_{i}]$, we have $0\in[-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j},\rho_{i}-\tau_{i}+{\eta}_{i-1}+{\eta}^{\prime}_{j}]$. This results in ${\eta}_{i-1}+{\eta}^{\prime}_{j}<\tau_{i}<+{\eta}_{i}+{\eta}^{\prime}_{j}$ which is equivalent to (13). Part 4) The sufficiency and necessity can be shown similar to Part 3. Part 1) We prove by contradiction that an equilibrium point of (7) should be either of these three types. Suppose $\bm{x}^{*}$ is an equilibrium point for (7) but is neither clean-cut, anticoordinator-driven, nor coordinator-driven. Now, based on the value of $x^{*}$, two cases may happen. Case 1) $x^{*}$ equals a threshold. Suppose that $x^{*}$ equals the threshold of one anticoordinating subpopulation, say $i$. From $\mathbf{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$ we have $0\in\mathcal{V}_{l}(\bm{x}^{*})$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, where for $l\neq i$, $\mathcal{V}_{l}(\bm{x}^{*})$ is a singleton and therefore to have $0\in\mathcal{V}_{l}(\bm{x}^{*})$, $x^{*}_{l}$ must equal $\rho_{l}$ for $l<i$ or $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, and $x^{*}_{l}$ must equal 00, otherwise, where $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ and $j=0$ if the set is empty. Therefore, $x^{*}-x^{*}_{i}={\eta}_{i-1}+{\eta}^{\prime}_{j}$. On the other hand, $x^{*}=\tau_{i}$ resulting in $x^{*}_{i}=\tau_{i}-{\eta}_{j}^{\prime}-{\eta}_{i-1}$ which is in turn equal to $a^{ij}_{i}$. This results in that $\bm{x}^{*}$ must be equivalent to $\mathbf{a}^{ij}$ which is a contradiction. Similarly, the case where $x^{*}$ equals the threshold of one coordinating subpopulation results in a contradiction. Therefore, $x^{*}$ cannot be equal to a threshold. Case 2) $x^{*}$ is not equal to any thresholds. In this case, $\bm{\mathcal{V}}(\bm{x}^{*})$ is a singleton and consequently, $\dot{\bm{x}}=\mathbf{0}$ which results in $\dot{x}={0}$. We showed that the equilibrium points of (11) should be either clean-cut, anticoordinator-driven, or coordinator-driven which is a contradiction. In view of Case 1 and Case 2, the assumption that an equilibrium point of (7) is neither clean-cut, anticoordinator-driven, nor coordinator-driven reaches a contradiction. ∎
VII-G A2
Lemma A2.
Consider the abstract dynamics (11). Between every two consecutive anticoordinator-driven and/or clean-cut equilibria, $q^{*}_{k-1}$ and $q^{*}_{k},k\in\{2,\ldots,\mathtt{Q}\}$, there exists one coordinator-driven equilibrium $q^{*}_{k-1,k}$ satisfying ${q}^{*}_{k-1}<q^{*}_{k-1,k}<{q}^{*}_{k}$.
Proof.
We prove by contradiction. Consider two consecutive clean-cut or anticoordinator-driven equilibrium points $q^{*}_{k-1}<q^{*}_{k}$. Assume on the contrary that there exist no coordinator-driven equilibrium points between $q^{*}_{k-1}$ and $q^{*}_{k}$. In view of the possible cases of $\mathcal{X}$ in (11), $\bar{f}(x)=\max\{y|y\in\mathcal{X}(x)\}$ will be negative for $x\in(q^{*}_{k-1},d^{+}_{k-1})$, where $d^{+}_{k-1}$ is lower bounded by $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, i.e., $d^{+}_{k-1}\geq\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, where $\bar{\tau}^{\prime}_{q_{k-1}}=\tau^{\prime}_{j+1}$, $\bar{\tau}_{q_{k-1}}=\tau_{i}$ when $q^{*}_{k-1}$ is clean-cut, i.e., $q^{*}_{k-1}=c^{ij}$ and $\bar{\tau}^{\prime}_{q_{k-1}}={\tau}^{\prime}_{j+1}$, $\bar{\tau}_{q_{k-1}}=\tau_{i-1}$ when $q^{*}_{k-1}$ is anticoordinator-driven, i.e., $q^{*}_{k-1}=a^{ij}$. Based on the value of $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}$, two cases may happen: Case 1) $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}=\bar{\tau}_{q_{k-1}}$ and consequently for a small enough $\varepsilon>0$, $\bar{f}(\bar{\tau}_{q_{k-1}}-\varepsilon)<0$. The sign of $\bar{f}(\bar{\tau}_{q_{k-1}}+\varepsilon)$ cannot be positive as $\bar{\tau}_{q_{k-1}}$ is the threshold of an anticoordinating subpopulation and considering the cases of $\mathcal{X}$ in (11), the value of $\dot{x}$ at $x=\tau_{i}+\varepsilon$ is not greater than its value at $x=\tau_{i}-\varepsilon$ for any $i\in[\mathsf{p}]$. Therefore, the sign of $\bar{f}$ does not change around $x=\bar{\tau}_{q_{k-1}}$. Actually, no matter how many thresholds of anticoordinating subpopulations (non benchmark subpopulations) are in the interval $(q^{*}_{k-1},q^{*}_{k})$, the sign of $\bar{f}$ will not change in the neighborhood of them. So, in this case, $\bar{f}(\bar{\tau}_{q_{k-1}}+\varepsilon)$ remains negative. Case 2) $\min\{\bar{\tau}^{\prime}_{q_{k-1}},\bar{\tau}_{q_{k-1}}\}=\bar{\tau}^{\prime}_{q_{k-1}}$ and consequently for a small enough $\varepsilon>0$, $\bar{f}({\bar{\tau}^{\prime}_{q_{k-1}}}-\varepsilon)<0$ and considering the second case of $\mathcal{X}$ in (11), the sign of $\bar{f}(\cdot)$ changes in a neighborhood of a coordinating subpopulation’s threshold only if $\bar{\tau}^{\prime}_{q_{k-1}}$ satisfies (14), i.e., $\bar{\tau}^{\prime}_{q_{k-1}}$ is a coordinator-driven equilibrium point. However, we assumed that it is not the case. Hence, the thresholds of coordinating subpopulations also do not impact the value of $d^{+}_{k-1}$. Putting the results of Case 1 and Case 2 together, we conclude that $d^{+}_{k-1}=q^{*}_{k}$, i.e., the sign of $\bar{f}(x)$ remains negative for $x\in(q^{*}_{k-1},q^{*}_{k})$. On the other hand, considering (12), (13), and the cases of $\mathcal{X}$ in (11), it can be concluded that $\underline{f}(x)=\min\{y|y\in\mathcal{X}(x)\}>0$ for $x\in(d^{-}_{k},q^{*}_{k})$, where $d^{-}_{k}$ is upper bounded by $\max\{\bar{\tau}^{\prime}_{q^{\prime}_{k}},\bar{\tau}_{q_{k}}\}$, where $\bar{\tau}^{\prime}_{q^{\prime}_{k}}={\tau}^{\prime}_{n}$, $\bar{\tau}_{q_{k}}=\tau_{m+1}$ when $q^{*}_{k}$ is clean-cut, i.e., $q^{*}_{k}=c^{mn}$ and similarly $\bar{\tau}^{\prime}_{q^{\prime}_{k}}=\tau^{\prime}_{n}$, $\bar{\tau}_{q_{k}}=\tau_{m+1}$ when $q^{*}_{k}$ is anticoordinator-driven, i.e., $q^{*}_{k}=a^{mn}$ As $x$ decreases, according to the cases of $\mathcal{X}$ in (11), the change of the sign of $\underline{f}(x)$ from positive to negative can only happen if the flow passes through the coordinating subpopulation’s threshold which equals a coordinator-driven equilibrium point, i.e., satisfies (14). However, we assumed it would not be the case. Hence, we conclude that the sign of $\underline{f}(x)$ is positive for $x\in(q^{*}_{k-1},q^{*}_{k})$. This is a contradiction as we already reached the conclusion that the sign of $\bar{f}(x)$ is negative for $x\in(q^{*}_{k-1},q^{*}_{k})$. Putting these arguments together, it is concluded that between two consecutive clean-cut or anticoordinator-driven equilibrium points, ${q}^{*}_{k-1},{q}^{*}_{k}$, there exists at least one coordinator-driven equilibrium point. Similarly, it can be shown that the number of coordinator-driven equilibrium points between two consecutive clean-cut/anti-coordinator-driven equilibria does not exceed one. We denote such equilibrium by ${q}^{*}_{k-1,k}$. Following similar arguments, it is straightforward to show that the leftmost and rightmost equilibrium points are clean-cut or anticoordinator-driven. Up to now, it has been shown that between two consecutive clean-cut or anticoordinator-driven equilibrium points of the abstract dynamics, $q^{*}_{k-1}$ and $q^{*}_{k}$, there exists a coordinator-driven equilibrium point, $q^{*}_{k-1,k}$ such that $q^{*}_{k-1}<q^{*}_{k-1,k}<q^{*}_{k}$. On the other hand, based Lemma 2, there is a one-to-one map between equilibrium points of the abstract dynamics and those of the continuous-time population dynamics. Therefore, for each two ordered clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics $\mathbf{q}^{*}_{k-1}$ and $\mathbf{q}^{*}_{k}$, there exists a coordinator-driven equilibrium point $\mathbf{q}^{*}_{k-1,k}$ such that $\smash{q^{*}_{k-1}<q^{*}_{k-1,k}<q^{*}_{k}}$. ∎
VII-H Proof of Theorem 2
As for Part 1, the instability of coordinator-driven equilibria is shown by providing an initial condition from which for a small $\epsilon$, there exists no $\delta$ to satisfy the definition of stability. The stability of clean-cut and anticoordinator-driven equilibria is first shown for the abstract dynamics using Theorem 1 in [51]. Then, the definition of asymptotic stability is used to show that of clean-cut and anticoordinator-driven equilibria for the continuous-time population dynamics. Part 2 of the theorem is shown by using the definition of the limit set.
Proof.
Part 1) Regarding the instability analysis of a coordinator-driven equilibrium point, say $\mathbf{q}^{*}_{k-1,k}=\mathbf{o}^{ij}$, consider the following initial condition $\bm{x}(0)=\mathbf{o}^{ij}+\delta^{*}\mathbf{e}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}$ for an arbitrarily small $\delta^{*}\in(0,\min\{\tau^{\prime}_{j+1},\tau_{i}\}-\tau^{\prime}_{j})$. The corresponding abstract state $x(0)$ will then belong to $(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. Pick $\epsilon$ to be equal to $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$. In view of (7), at such $x(0)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton. Moreover, considering the structure of $\mathbf{o}^{ij}$, at $x(0)$ we have $\dot{x}_{l}=0$ for $l\in\{1,\ldots,\mathsf{p}+\mathsf{p}^{\prime}-j,\mathsf{p}+\mathsf{p}^{\prime}-j+2,\ldots,\mathsf{p}+\mathsf{p}^{\prime}\}$. Indeed, the equation $\dot{x}_{l}=0$ holds true as long as $x\in(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. On the other hand, $x(0)>\tau^{\prime}_{j}$ and in view of (14), we have $\dot{x}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\rho^{\prime}_{j}-x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)>0$, for all $t\leq t^{\prime}$, where $t^{\prime}=\inf\{t\geq 0|x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}\}$, where $\alpha_{i,j}=\min\{\tau^{\prime}_{j+1},\tau_{i}\}-({\eta}_{i}+{\eta}_{j-1})$. If $\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}=\alpha_{i,j}$, then $x(t^{\prime})=\min\{\tau^{\prime}_{j+1},\tau_{i}\}$. Otherwise, $x(t^{\prime})={\eta}_{j}+{\eta}_{i}$. In either case, we observe that $|x(t^{\prime})-\tau_{j}^{\prime}|$ is greater than the selected value for $\epsilon$, i.e., $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$, and the value of $\delta^{*}$ does not impact the value of $x(t^{\prime})$. This indicates instability of ${o}^{ij}$ and in turn $\mathbf{o}^{ij}$. The stability properties of the other two types of equilibrium points, are first shown for the abstract dynamics. Let $V_{k}(x):\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}=\frac{1}{2}(x-q^{*}_{k})^{2}$, where $q^{*}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}$ if $q^{*}_{k}$ is clean-cut, i.e., $q^{*}_{k}=c^{ij}$ and $q^{*}_{k}=\tau_{i}$ if it is anticoordinator-driven, i.e., $q^{*}_{k}=a^{ij}$. The set-valued derivative of $V_{k}(x)$ w.r.t (11) will be $\mathcal{D}(x)=\{(x-q^{*}_{k})\nu|\nu\in\mathcal{X}(x)\}$ [51]. We also have $\bar{f}(x)=\max\{y|y\in\mathcal{X}(x)\}<0$ for $x\in(q^{*}_{k},q^{*}_{k,k+1})$ and $\underline{f}(x)=\min\{y|y\in\mathcal{X}(x)\}>0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k})$ (refer to the proof of A2). Hence, $\max\mathcal{D}(x)<0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})/\{q^{*}_{k}\}$. According to [51, Theorem 1], $q^{*}_{k}$ is a strongly asymptotically stable equilibrium point for (11). Also the basin of attraction of $q^{*}_{k}$, will be $(q^{*}_{k-1,k},q^{*}_{k,k+1})$. As for $q^{*}_{1}$ (resp. $q^{*}_{\mathtt{Q}}$), we have $\bar{f}(0)>0$ (resp. $\bar{f}(1)<0$) and this case can be handled similarly which results in $\mathcal{A}(q^{*}_{1})=[0,q^{*}_{1,2})$ (resp. $\mathcal{A}(q^{*}_{\mathtt{Q}})=(q^{*}_{\mathtt{Q}-1,\mathtt{Q}},1]$). Now, we show the asymptotic stability of $\mathbf{q}^{*}_{k}$ when it is an anticoordinator-driven equilibrium point, i.e., $\mathbf{q}^{*}_{k}=\mathbf{a}^{ij}$. The case where $\mathbf{q}^{*}_{k}$ is clean-cut can be handled similarly. We show that for every $\varepsilon>0$, there exists a $\delta>0$ such that if $|\bm{x}(0)-\mathbf{a}^{ij}|<\delta$, then $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$ for all $t\geq 0$, and $|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Consider a small enough $\epsilon>0$, such that at an $\epsilon$-neighborhood of $a^{ij}$, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ or for $l$ satisfying both $\mathsf{p}<l$ and $0<\mathsf{p}+\mathsf{p}^{\prime}+1-l\leq j$. For the remaining $l$’s, $\dot{x}_{l}=-x_{l}$. From the stability of the abstract dynamics, the existence of $\delta_{x}>0$ follows such that $|x(0)-a^{ij}|<\delta_{x}$ results in $|x(t)-a^{ij}|<\epsilon$ for $t\geq 0$. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, the final value of $x_{l}$ is $a^{ij}_{l}$ and consequently $x_{l}$ will approach $a^{ij}_{l}$ exponentially, i.e.,
$|x_{l}(t)-{a}^{ij}_{l}|\leq(x_{l}(0)-a^{ij}_{l})e^{-t},$ (18)
for $t\geq 0$. On the other hand, we know that for any $y,z\in\mathbb{R}$, $|y+z|\leq|y|+|z|$. Hence, in view of $|x_{i}-a^{ij}_{i}|=$ $|(x-a^{ij})-\Sigma_{l\neq i}(x_{l}-a^{ij}_{l})|$ , we have
$|x_{i}-a^{ij}_{i}|\leq|(x-a^{ij})|+|\Sigma_{l\neq i}(x_{l}-a^{ij}_{l})|.$ (19)
Consequently, for $t\geq 0$
$|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<(|x(t)-a^{ij}|+|\Sigma_{l\neq i}(x_{l}-a^{ij}_{l})|)^{2}+\Sigma_{l\neq i}|x_{l}-a^{ij}_{l}|^{2}.$
On the other hand, starting from $\delta_{x}$-neighborhood of $a^{ij}$, in view of (18), we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<\big{(}(\mathsf{p}+\mathsf{p}^{\prime})\max_{l}|x_{l}(0)-a^{ij}_{l}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t}\big{)}^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$ and, in turn, $|\bm{x}(t)-\mathbf{a}^{ij}|$ $<\max_{l}|x_{l}(0)-a^{ij}_{l}|\big{(}\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-2t}$$\big{)}^{0.5}.$ If $\delta$ is less than $\delta_{x}/(\mathsf{p}+\mathsf{p}^{\prime})$, then in view of the relation between Euclidean norm and Norm-1, $\sum_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}|x_{l}(0)-a_{l}^{ij}|$ will be less than $\delta_{x}/\sqrt{\mathsf{p}+\mathsf{p}^{\prime}}$. and, in turn, $|x(0)-a^{ij}|<\delta_{x}$. Hence, $\mathbf{a}^{ij}$ is stable because by choosing $\delta=\min\{\frac{\delta_{x}}{\mathsf{p}+\mathsf{p}^{\prime}},\frac{\varepsilon}{\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1))^{2}+(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}\}},$ we have $|x(t)-a^{ij}|<\epsilon$ for all $t\geq 0$, and consequently $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$. As for asymptotic stability, we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<(|x(t)-a^{ij}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$, for $t\geq 0$. Since each term in the right hand side approaches zero, we have $\lim|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Now, we obtain the basin of attraction of $\mathbf{q}^{*}_{k}$. (i) It is straightforward to show that if $x(0)$ falls in the basin of attraction $q^{*}_{k}$, the abstract dynamics will converge to an arbitrary small neighborhood of $q^{*}_{k}$ in a finite time. Now, depending on the types of $q^{*}_{k}$, two cases may happen: Case 1) $q^{*}_{k}=c^{ij}$, i.e., $q^{*}_{k}$ is clean-cut, then let $t=t_{0}$ be the time instant at which the abstract state $x$ enters a small enough $\epsilon$-neighborhood of $q^{*}_{k}$, such that for $l\leq i$ or $l\geq\max\{\mathsf{p}+\mathsf{p}^{\prime}-j+1,\mathsf{p}+1\}$, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining $l$’s, we have $\dot{x}_{l}=-x_{l}$. Then for $t>t_{0}$, the asymptotic value of $x_{l}$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$ is the same as the $l$-th entry of $\mathbf{q}^{*}_{k}$. Moreover, each $x_{l}$ approaches $c^{ij}_{l}$ exponentially. As a result, for $t>t_{0}$ we have $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<C\exp(-{(}t{-t_{0}}))$, for some constant $C$. Case 2) $q^{*}_{k}=a^{ij}$, i.e., $q^{*}_{k}$ is anticoordinator-driven. Then, a time moment similar to $t_{0}$ defined in Case 1 exists such that for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $|x_{l}(t)-a^{ij}_{l}|<C\exp(-{(}t{-t_{0}}))$, for some constant $C$ and $t\geq t_{0}$. Therefore, in view of (19), it is straightforward to show that for any arbitrary small $\epsilon_{1}$, there exists some finite time $t_{1}+t_{0}$, such that $|x_{i}(t)-{q}^{*}_{k_{i}}|<\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}$ for $t>t_{1}+t_{0}$ resulting in $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<(\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}^{2}$. Having this and (i) we conclude that in both cases 1 and 2, for any arbitrary $\varepsilon$-neighborhood of $\mathbf{q}^{*}_{k}$, time $T>0$ can be found such that for all $t\geq T$, $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<\varepsilon$. As a result, the basin of attraction of point $\bm{q}^{*}_{k}$ equals $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})\}$. The basins of attraction of $\mathbf{q}^{*}_{1}$ and $\mathbf{q}^{*}_{\mathtt{Q}}$ can be obtained similarly. Part 2) When $x(0)=q^{*}_{k-1,k}$, where $q^{*}_{k-1,k}=o^{ij}$, two cases may happen. Case 1) The abstract state $x$ remains at $q^{*}_{k-1,k}$. Then for $l\neq j$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}$ is $\rho_{l}-x_{l}$ for $l\leq i$, or for $l>\mathsf{p}$ and $\mathsf{p}+\mathsf{p}^{\prime}+1-j<l$, and $\dot{x}_{l}$ is $-x_{l}$ for the remaining $l$’s. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{j\}$, the final value of $x_{l}$ is $o^{ij}_{l}$ and consequently $x_{l}$ will approach $o^{ij}_{l}$ exponentially. Then for any arbitrary small $\varepsilon$, in a finite time $T$ we have $|x_{l}(t)-o^{ij}_{l}|<\varepsilon$ for $t\geq T$ and subsequently $|x_{j}(t)-o^{ij}_{j}|<|x(t)-o^{ij}|+\Sigma_{l\neq j}|(x_{l}(t)-o^{ij}_{l})|\leq(\mathsf{p}+\mathsf{p}^{\prime}-1)\varepsilon$ resulting in $|\bm{x}(t)-\mathbf{o}^{ij}|\leq\varepsilon\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}$ for $t\geq T$. Hence, $\mathbf{o}^{ij}$ is the limit point. Case 2) The abstract state $x$ leaves $q^{*}_{k-1,k}$. Then, the provided reasoning for the case $x\neq q^{*}_{k-1,k}$ will be applicable resulting in $\mathbf{q}^{*}_{k-1}$ or $\mathbf{q}^{*}_{k}$ to be the limit point. Overall, the limit set of $\mathbf{q}^{*}_{k-1,k}$ is $\{\mathbf{q}^{*}_{k-1},\mathbf{q}^{*}_{k-1,k},\mathbf{q}^{*}_{k}\}$. ∎
VII-I Proof of Proposition 3
Proof.
Based on Theorem 2, the set $\mathbf{Q}^{c}\cup\mathbf{Q}^{a}\cup\mathbf{Q}^{o}$ contains the limit sets of all $\bm{x}$ in $\bm{\mathcal{X}}_{ss}$. To find the corresponding Birkhoff center, we need to find the closure of the set of recurrent points. Based on Definition 3, the set of recurrent points is also equal $\mathbf{Q}^{c}\cup\mathbf{Q}^{a}\cup\mathbf{Q}^{o}$. As a results, the Birkhoff center of the population dynamics will be equal to $\mathbf{Q}^{c}\cup\mathbf{Q}^{a}\cup\mathbf{Q}^{o}$. ∎
VII-J Proof of Theorem 3
Proof.
(i) Based on 1, the sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle$ which evolves according to discrete population dynamics (3), is a realization of population dynamics Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ with transition probabilities formulated in (6). (ii) Based on Lemma 1, the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the good upper semicontinuous differential inclusion (7). (iii) The transition probabilities of the population dynamics Markov chain (6) are homogeneous and the state space over which the Markov chain is defined is finite for each $\mathsf{N}$. This results in the existence of invariant probability measures ${\mu}^{\frac{1}{\mathsf{N}}}$ for Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ [54]. (iv) 3 specifies the Birkhoff centers of (7). Consider the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence. Theorem 1 and (i)-(iv) together complete the proof. ∎
VII-K Proof of 4
Proof.
Assume on the contrary that there exists some $N>0$, such that for every population size $\mathsf{N}\in\mathcal{N}_{\geq N}$, where $\mathcal{N}_{\geq N}=\{\mathsf{N}|\mathsf{N}\in\mathcal{N},\mathsf{N}\geq N\}$, the evolution of the proportion of $\mathtt{A}$-players admits some perpetual fluctuation set ${\mathcal{Y}}^{\mathsf{N}}\subset(0,1)\cap\frac{1}{\mathsf{N}}\mathbb{Z}$ of amplitude $y^{\mathsf{N}}$, where $y^{\mathsf{N}}\geq L$ for some positive value $L$,
$\exists L>0\forall\mathsf{N}\in\mathcal{N}_{\geq N}\qquad y^{\mathsf{N}}\geq L$ (20)
and that for every $K\in(0,L]$ there exists some small enough $\delta>0$ such that in the long run the proportion of $\mathtt{A}$-players visits every connected subset of amplitude $K$ of ${\mathcal{Y}}^{\mathsf{N}}$ (for some $K>0$) with probability greater than $\delta$:
$\forall K>0\exists\delta>0\forall$ $\mathcal{Z}^{\mathsf{N}}\subset[\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]\text{ with }|\mathcal{Z}^{\mathsf{N}}|=K$
[…]
where $\mu_{x}^{\frac{1}{\mathsf{N}}}$ is defined as $\mu^{\frac{1}{\mathsf{N}}}\circ\bm{f}^{-1}$, where $\bm{f}:\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}\to\mathbb{R},$ and $\bm{f}(\bm{x})=\sum_{p=1}^{\mathsf{p}}x_{p}.$ Let $\mathtt{n}$ denote the total number of abstract equilibrium points, and let $q_{p}$, for $p\in[\mathtt{n}]$, denote the $p^{\text{th}}$ equilibrium point of the abstract dynamics. Given Theorem 3, it follows that for every vanishing sequence $\langle\frac{1}{n}\rangle=\frac{1}{n_{0}},\frac{1}{n_{1}},\ldots$ where $n_{i}\in\mathcal{N}$ for $i\in\mathbb{Z}_{\geq 0}$ $\lim_{i\to\infty}\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))=1,$ for a small enough positive $\epsilon<L/\mathtt{n}$. Equivalently, for every $\delta>0$, there exists $i^{*}>0$ such that for all $i>i^{*}$, $\smash{\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))>1-\delta}$ and, in turn, $\smash{\mu_{x}^{\frac{1}{n_{i}}}([\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))<\delta}$. On the other hand, in view of (20), $[\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2)$ contains a connected interval of amplitude at least $(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$, This contradicts (21) for $K=(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$. ∎
VII-L A3
Lemma A3.
The population dynamics governed by update rules (1) and (2) and the activation sequence are equivalent to the dynamics defined in (3).
Proof.
As the update mechanism is asynchronous, the population states at two consecutive time indices $\bm{x}^{\mathsf{N}}(k)$ and $\bm{x}^{\mathsf{N}}(k+1)$ can differ by at most ±$\pm\frac{1}{\mathsf{N}}$. Which element of the population state $\bm{x}^{\mathsf{N}}(k+1)$ may differ from that of $\bm{x}^{\mathsf{N}}(k)$ is determined by the random variable ${P}_{k}$ which is the active subpopulation at time index $k$, and can take values in $[\mathsf{p}+\mathsf{p}^{\prime}]$. The probability of ${P}_{k}=p$ equals $\rho_{p}$. The amount of change in the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ depends on the current strategy of the active agent and her preferred strategy. The former is captured by the random variable ${S}_{k}$ taking values $1$ and $2$, and the latter is determined by the function $s^{*}({P}_{k},\bm{x}^{\mathsf{N}}(k),{S_{k}})$ which is obtained from (1) and (2). The probability of ${S}_{k}=1$ (resp. ${S}_{k}=2$) given the active subpopulation is $p$ is $x^{\mathsf{N}}_{p}/\rho_{p}$ (resp. $1-x^{\mathsf{N}}_{p}/\rho_{p}$). If the current strategy and the preferred strategy are equal, the population states at time indices $k$ and $k+1$ are the same, and if the current strategy is $\mathtt{B}$ and the preferred one is $\mathtt{A}$, the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ increases by $\frac{1}{\mathsf{N}}$, and the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ decreases by $\frac{1}{\mathsf{N}}$ otherwise. This update rule is equivalent to (3). ∎
Corollary A1.
Assume that an active agent $i$ chooses her strategy uniformly at random when $x_{-i}^{\mathsf{N}}(k)=\tau(i)$. Then, Lemma 1 remains true.
Proof.
It is easy to show that Conditions 1, 2, and 4 in Definition 4 are satisfied. Now, we investigate Condition 3. Function $s^{*}(\cdot,\cdot,\cdot)$ changes to the random variable $S^{*}$ with the following distribution
$\mathbb{P}[$ $S^{*}(p,\bm{x}^{\mathsf{N}},s)=1\mid\bm{x}^{\mathsf{N}}]=$
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{or }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 1&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\end{cases}$
$\mathbb{P}[$ $S^{*}(p,\bm{x}^{\mathsf{N}},s)=2\mid\bm{x}^{\mathsf{N}}]$
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ 1&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1.\\ \end{cases}$
Following the steps taken in the proof of Lemma 1, the expected change in the population state $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ can be obtained, where for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ satisfying $\nexists\tau_{p}(\bm{1}^{\top}\bm{x}\in[\tau_{p},\tau_{p}+\frac{1}{\mathsf{N}}])$ is independent of the tie-breaking rule. Out of those states satisfying $\exists\tau_{p}(\bm{1}^{\top}\bm{x}\in[\tau_{p},\tau_{p}+\frac{1}{\mathsf{N}}])$, we investigate the state $\bm{x}$ with $\bm{1}^{\top}\bm{x}=\tau_{p}$ for some $p$; the other values of $\bm{x}$ can be handled similarly. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, which corresponds to subpopulation $p$, at $\bm{x}$ with $\bm{1}^{\top}\bm{x}=\tau_{p}$ equals $\mathsf{N}\left(\frac{1}{\mathsf{N}}(\rho_{p}-x_{p})\frac{1}{2}-\frac{1}{\mathsf{N}}x_{p}\times\big{(}0\big{)}\right)=\frac{\rho_{p}}{2}-\frac{x_{p}}{2},$ if $p$ corresponds to an anticoordinating subpopulation and $\mathsf{N}\left(\frac{1}{\mathsf{N}}(\rho_{p}-x_{p})\frac{1}{2}-\frac{1}{\mathsf{N}}x_{p}\big{(}1\big{)}\right)=\frac{\rho_{p}}{2}-\frac{3}{2}x_{p},$ otherwise. In both cases the expected change in subpopulation $p$ belongs to $\mathcal{V}_{p}(\bm{x})=[-x_{p},\rho_{p}-x_{p}]$ (equation (7)). For elements $l\neq p$ of $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, the expected change is equal to the unique member of the $\mathcal{V}_{l}(\bm{x})$. This implies that equation (5) is satisfied with $\bm{y}=\bm{x}$. Hence, Condition 3 holds, and this completes the proof. ∎